### v1

# A Finer View of the Parameterized Landscape of Labeled Graph Contractions

Abstract
We study the Labeled Contractibility problem, where the input consists of two vertex-labeled graphs $G$ and $H$, and the goal is to determine whether $H$ can be obtained from $G$ via a sequence of edge contractions.
Lafond and Marchand [WADS 2025] initiated the parameterized complexity study of this problem, showing it to be ${\mathsf{W}}[1]$-hard when parameterized by the number $k$ of allowed contractions. They also proved that the problem is fixed-parameter tractable when parameterized by the tree-width tw of $G$, via an application of Courcelle’s theorem resulting in a non-constructive algorithm.
In this work, we present a constructive fixed-parameter algorithm for Labeled Contractibility with running time $2^{\mathcal{O}(\textsf{tw}^{2})}\cdot|V(G)|^{\mathcal{O}(1)}$. We also prove that unless the Exponential Time Hypothesis (ETH) fails, it does not admit an algorithm running in time $2^{o(\textsf{tw}^{2})}\cdot|V(G)|^{\mathcal{O}(1)}$. This result adds Labeled Contractibility to a small list of problems that admit such a lower bound and matching algorithm.
We further strengthen existing hardness results by showing that the problem remains $\mathsf{NP}$-complete even when both input graphs have bounded maximum degree. We also investigate parameterizations by $(k+\delta(G))$ where $\delta(G)$ denotes the degeneracy of $G$, and rule out the existence of subexponential-time algorithms. This answers question raised in Lafond and Marchand [WADS 2025]. We additionally provide an improved $\mathsf{FPT}$ algorithm with better dependence on $(k+\delta(G))$ than previously known. Finally, we analyze a brute-force algorithm for Labeled Contractibility with running time $|V(H)|^{\mathcal{O}(|V(G)|)}$, and show that this running time is optimal under ETH.
Keywords: Labeled Contraction, ETH Lower Bound, Treewidth
1 Introduction
Graph modification problems are a foundational topic in algorithmic graph theory. A general formulation of the $\mathcal{F}$-M-Modification problem asks: given a graph $G$ and a fixed set M of allowed modification operations (such as vertex deletion, edge addition/deletion, or edge contraction), can $G$ be transformed into a graph in class $\mathcal{F}$ using at most $k$ operations from M.
It is well known that the $\mathcal{F}$-M-Modification problem is $\mathsf{NP}$-hard for many natural graph classes $\mathcal{F}$ when M consists of vertex or edge deletions [54, 42]. Hardness results for edge contraction were first shown by Watanabe et al. [53] and Asano and Hirata [4]. It is noteworthy that while the problem becomes trivially solvable when M includes only vertex or edge deletions and $\mathcal{F}$ is a singleton, this is not the case for edge contractions. In fact, Brouwer and Veldman [9] established that the problem remains $\mathsf{NP}$-complete even when $\mathcal{F}$ is a singleton containing a small graph, such as a cycle or a four-vertex path. This difference highlights the increased algorithmic complexity introduced by edge contractions compared to vertex or edge deletions.
Graph modification problems have played a central role in the development of parameterized complexity and the design of key algorithmic techniques. A representative list of relevant works includes [7, 8, 11, 12, 13, 14, 17, 23, 24, 25, 27, 29]. For a comprehensive overview, the survey by Crespelle et al. [18] provides an in-depth discussion of modification problems involving edge deletion, editing, and completion.
A series of recent papers have studied the parameterized complexity of $\mathcal{F}$-Contraction for various graph classes $\mathcal{F}$, including paths and trees [35], generalizations and restrictions of trees [1, 2], cactus graphs [40], bipartite graphs [34, 36], planar graphs [33], grids [50], cliques [10], split graphs [3], chordal graphs [44], bicliques [47], and degree-constrained graph classes [5, 32, 51].
Recently, Lafond and Marchand [41] initiated the parameterized study of edge contraction problems in uniquely labeled graphs, where each vertex has a different label. They introduced the following problem:
Labelled Contractibility Input: Two vertex-labeled graphs $G$ and $H$ with $V(H)\subseteq V(G)$. Question: Is $H$ a labeled contraction of $G$?
See Definition 2.1 in Section 2 for formal definitions.
They also introduced a generalization, motivated by applications in phylogenetic networks [46], which are rooted acyclic directed graphs with labeled leaves representing evolutionary histories:
Maximum Common Labeled Contraction Input: Two vertex-labeled graphs $G$ and $H$, and an integer $k$. Question: Do there exist labeled contraction sequences $S_{1}$ and $S_{2}$ such that $G/S_{1}=H/S_{2}$ and $|S_{1}|+|S_{2}|\leq k$?
This second problem is closely related to the first. As shown in [41], $H$ is a labeled contraction of $G$ if and only if there exists a common contraction of both $G$ and $H$ to a labeled graph $H^{\prime}$ of size at least $|V(H)|$. This equivalence allows complexity results to be transferred between the two problems. The same work provides additional applications and results.
Lafond and Marchand [41] showed that Labeled Contractibility is ${\mathsf{W}}[1]$-hard when parameterized by the natural parameter $k=|V(G)\setminus V(H)|$, the number of allowed edge contractions. This motivates the exploration of structural parameters. They established that the problem admits a fixed-parameter tractable ($\mathsf{FPT}$) algorithm when parameterized by the treewidth tw of the input graph $G$ by invoking a variation of Courcelle’s theorem [26, Theorem 11.73]. However, this approach yields a non-elementary dependence on the parameter.
Our first contribution is a constructive dynamic programming algorithm with a running time that is singly exponential in $\textsf{tw}^{2}$. We complement this with a matching lower bound under the Exponential Time Hypothesis (ETH):
Theorem 1.
Labeled Contractibility
  • admits an algorithm running in time $2^{\mathcal{O}(\textsf{tw}^{2})}\cdot|V(G)|^{\mathcal{O}(1)}$; but
  • does not admit an algorithm running in time $2^{o(\textsf{tw}^{2})}\cdot|V(G)|^{\mathcal{O}(1)}$, unless the ETH fails.
We note that this lower bound also holds when parameterized by the pathwidth, which is a larger parameter. Most known $\mathsf{FPT}$-algorithms that use dynamic programming on tree decompositions have running times, and matching lower bounds, of the form $2^{\mathcal{O}(\textsf{tw})}\cdot n^{\mathcal{O}(1)}$ [19, Chapters 7,13] or $2^{\mathcal{O}(\textsf{tw}\log\textsf{tw})}\cdot n^{\mathcal{O}(1)}$ [43, 21]. For a relatively rare set of problems exhibits a double-exponential dependence on treewidth, see  [30, 52, 16] and references within. To our knowledge, the only problems with single-exponential-but-polynomial dependencies on pathwidth are mentioned in [20, 48, 49]. Similar results parameterized by vertex cover number can be found in [3, 31, 6, 15].
Lafond and Marchand [41] observed that smaller parameters such as degeneracy do not yield tractable results. Specifically, it is known that the Maximum Common Labeled Contraction problem is $\mathsf{NP}$-hard even when both input graphs have constant maximum degree (and thus constant degeneracy), and that Labeled Contractibility is $\mathsf{NP}$-hard for graphs of bounded degeneracy [45, Theorem 8]. We strengthen this latter result by proving hardness under an even stricter constraint:
Theorem 2.
Labeled Contractibility is $\mathsf{NP}$-hard even when both $G$ and $H$ have bounded maximum degree.
These results suggest that tractability is unlikely when parameterizing solely by the solution size $k$ or the degeneracy $\delta(G)$. Nevertheless, Lafond and Marchand [41] showed the problem becomes tractable for the combined parameter $(k+\delta(G))$, yielding an algorithm with a running time of $(\delta(G)+2k)^{k}\cdot n^{\mathcal{O}(1)}$. They asked whether a subexponential-time algorithm exists for this parameterization, a question we resolve negatively:
Theorem 3.
Labeled Contractibility does not admit an algorithm running in time $2^{o(|V(G)|+|E(G)|)}$ even when both $G$ and $H$ have bounded degeneracy, unless the ETH fails.
This theorem rules out the possibility of a subexponential-time algorithm parameterized by $(k+\delta(G))$ under the assumption of the ETH. We also provide the following algorithm, which improves upon previous work:
Theorem 4.
Labeled Contractibility can be solved in time $(\delta(H)+1)^{k}\cdot|V(G)|^{\mathcal{O}(1)}$, where $k$ is the solution size and $\delta(H)$ is the degeneracy of $H$.
This algorithm improves upon the $(\delta(H)+k)^{k}\cdot n^{\mathcal{O}(1)}$ algorithm from [41, Section 3.3]. The previous result relied on the bound $\delta(H)\leq\delta(G)+k$. We refine this with a tighter upper bound: $\delta(H)\leq\delta(G)\cdot\frac{|V(G)|}{|V(G)|-k}$. For cases where $|V(G)|\geq(1+\epsilon)\cdot k$, this bound yields $\delta(H)\leq\delta(G)\cdot c_{\epsilon}$, where $\epsilon$ and $c_{\epsilon}$ are absolute constants. This substantially improves the previous analysis, leading to a more efficient algorithm under these conditions.
Finally, as our last contribution, we analyze a brute-force algorithm for Labeled Contractibility and establish its optimality under the ETH. Notably, our analysis also applies to the more general Maximum Common Labeled Contraction problem.
Theorem 5.
Labeled Contractibility
  • admits an algorithm running in time $|V(H)|^{\mathcal{O}(|V(G)|)}$; but
  • does not admit an algorithm running in time $|V(H)|^{o(|V(G)|)}$, unless the ETH fails.
Organization. We adopt standard notation from graph theory and parameterized complexity, as outlined in Section 2. In Section 3, we present a dynamic programming algorithm based on tree decompositions and establish a matching lower bound. Specifically, we provide a parameter-preserving reduction from the Sub-Cubic Partitioned Vertex Cover problem [3], which proves Theorem 1. Section 4 proves Theorem 2 by a reduction from Positive-Not-All-Equal-SAT, showing $\mathsf{NP}$-hardness even for graphs of bounded maximum degree. Section 5 is dedicated to a conditional lower bound under the ETH. Using a reduction from 1-in-3-SAT, we prove this result even when both input graphs have bounded degeneracy. We also present an improved $\mathsf{FPT}$ algorithm for the combined parameter $(k+\delta(G))$. These results together establish Theorems 3 and 4. Section 6 analyzes a brute-force algorithm and establishes its optimality under the ETH. This is achieved via a reduction from the Cross Matching problem [28] and establishes Theorem 5. Finally, Section 7 concludes with open problems and conjectures.
2 Preliminaries
For a positive integer $n$, we denote $n]=\{1,2,\ldots,n\}$. Let $G$ be a graph. We denote by $V(G)$ and $E(G)$ the vertex set and edge set of $G$, respectively. Two vertices $u,v\in V(G)$ are said to be adjacent if $(u,v)\in E(G)$. For a vertex $u\in V(G)$, let $N_{G}(u)$ denote the (open) neighborhood of $u$, i.e., the set of all vertices adjacent to $u$ in $G$, and define the closed neighborhood as $N_{G}[u]=N_{G}(u)\cup\{u\}$. We omit the subscript if the graph is clear from the context. Given two disjoint subsets $X,Y\subseteq V(G)$, we say that $X$ and $Y$ are adjacent if there exists an edge in $G$ with one endpoint in $X$ and the other in $Y$. For any $X\subseteq V(G)$, the subgraph of $G$ induced by $X$ is denoted by $G[X]$. We write GXG-X to denote the induced subgraph $G[V(G)\setminus X]$. The maximum degree of $G$, denoted $\Delta(G)$, is defined as $\max_{u\in V(G)}|N_{G}(u)|$. The degeneracy of $G$, denoted $\delta(G)$, is the smallest integer $d$ such that every subgraph of $G$ contains a vertex of degree at most $d$.
To cope up with the hardness of $\mathsf{NP}$-hard problems, these problems are studied through the framework of parameterized complexity. This approach enables a finer analysis of computational problems by introducing a parameter $\ell$ (e.g., solution-size, treewidth, vertex cover number), which may be significantly smaller than the input size. A problem is fixed-parameter tractable ($\mathsf{FPT}$) if it admits an algorithm running in time $f(\ell)\cdot|I|^{\mathcal{O}(1)}$ for some computable function $f$ and instance size $|I|$. Problems that are hard for the class ${\mathsf{W}}[1]$ are believed not to admit such algorithms under standard assumptions, and those that remain $\mathsf{NP}$-hard even for fixed parameter values are called para-$\mathsf{NP}$-hard. We refer readers to book [19] for other related terms.
The Exponential Time Hypothesis (ETH) of Impagliazzo and Paturi [37] implies that the 3-Sat problem on $n$ variables cannot be solved in time $2^{o(n)}$. The Sparsification Lemma of Impagliazzo et al. [38] implies that if the ETH holds, then there is no algorithm solving a 3-Sat formula with $n$ variables and $m$ clauses in time $2^{o(n+m)}$.
Throughout this work, we consider uniquely labeled graphs. Following the convention in [41], we refer to vertices directly by their labels rather than using explicit labeling functions. Two labeled graphs $G$ and $H$ are considered equal, denoted $G=H$, if and only if $V(G)=V(H)$ and $E(G)=E(H)$. This is distinct from the standard notion of graph isomorphism, where a bijection $\sigma:V(G^{\prime})\to V(H^{\prime})$ exists such that $(u,v)\in E(G^{\prime})$ if and only if $(\sigma(u),\sigma(v))\in E(H^{\prime})$. In our setting, we assume that the labels of $H$ are a subset of those of $G$.
Definition 2.1.
Let $G$ be a labeled graph and $(u,v)\in E(G)$. The labeled contraction of the edge $(u,v)$ is an operation that transforms $G$ by:
  • For every vertex $w\in N(v)\setminus N[u]$, adding the edge $(u,w)$ to $G$.
  • Removing the vertex $v$ and all incident edges.
The resulting graph is denoted by $G/(u,v)$.
Note that in general, $G/(u,v)\neq G/(v,u)$, as the vertex that is retained after the contraction (and thus the label) is different. The two resulting graphs are, however, isomorphic under the standard unlabeled notion. For simplicity, in the remainder of this paper, we assume all contractions are labeled and refer to them simply as contractions.
A labeled contraction sequence $S$ on a graph $G$ is a sequence of vertex pairs $S=((u_{1},v_{1}),\ldots,(u_{k},v_{k}))$ such that for each $i\in\{0,\ldots,k-1\}$, if $G^{i}$ is the graph obtained after the first $i$ contractions (with $G^{0}:=G$), then $(u_{i+1},v_{i+1})\in E(G^{i})$ and $G^{i+1}=G^{i}/(u_{i+1},v_{i+1})$. The graph obtained after the full sequence is denoted by $G/S$. If $S$ is not a valid contraction sequence on $G$, then $G/S$ is undefined. We say that a graph $H$ is a labeled contraction of $G$ if there exists a contraction sequence $S$ such that $G/S=H$. A graph $M$ is a common labeled contraction of graphs $G$ and $H$ if it is a labeled contraction of both. A maximum common labeled contraction of $G$ and $H$ is a common labeled contraction with the largest possible number of vertices.
Consider two graphs $H$ and $G$ such that $V(H)\subseteq V(G)$.
Definition 2.2.
A witness structure of $G$ into $H$ is a partition $\mathcal{W}=\{W_{1},\ldots,W_{|V(H)|}\}$ of $V(G)$ into non-empty sets satisfying:
  • For each $W_{i}\in\mathcal{W}$, the induced subgraph $G[W_{i}]$ is connected.
  • Each $W_{i}$ contains exactly one vertex from $V(H)$, called its representative vertex.
  • For any distinct $u,v\in V(H)$, the edge $(u,v)\in E(H)$ if and only if the sets in $\mathcal{W}$ containing $u$ and $v$ are adjacent in $G$.
There is a natural bijection between the sets in $\mathcal{W}$ and the vertices of $H$ via their representatives. It is known that $H$ is a labeled contraction of $G$ if and only if a witness structure of $G$ into $H$ exists (see Observation 2 in [41]). Furthermore, contractions within a single witness set can be performed in any order as long as the representative vertex is preserved (Observation 3 in the same work). For any vertex $u\in V(G)$, we denote by $\mathcal{W}(u)\in\mathcal{W}$ the unique set that contains $u$.
3 Parameterized by Treewidth
In this section, we present the proof of Theorem 1 which we restate.
See 1
In the first subsection, we describe the algorithmic result by presenting a dynamic programming algorithm on tree decomposition. In the second subsection, we establish a matching conditional lower bound.
3.1 Algorithmic Result
In this subsection we will present a dynamic programming (DP) based algorithm for the Labeled Contraction problem when parameterized by treewidth of the input graph. For the sake of completeness, we start with the standard definitions.
Definition 3.1 ([19]).
A tree decomposition of a graph $G$ is a pair $(\mathcal{T},\{X_{t}\}_{t\in V(\mathcal{T})})$, where $\mathcal{T}$ is a rooted tree and each node $t\in V(\mathcal{T})$ is associated with a set $X_{t}\subseteq V(G)$, referred to as a bag, satisfying the following conditions:
  • For every vertex $v\in V(G)$, the set $\{t\in V(\mathcal{T})\mid v\in X_{t}\}$ is nonempty and induces a connected subtree of $\mathcal{T}$;
  • For every edge $\{u,v\}\in E(G)$, there exists a node $t\in V(\mathcal{T})$ such that $\{u,v\}\subseteq X_{t}$.
The width of a tree decomposition is $\max_{t\in V(\mathcal{T})}|X_{t}|-1$. The treewidth of $G$, denoted $\textsf{tw}(G)$, is the minimum width over all valid tree decompositions of $G$.
Definition 3.2 ([19]).
A nice tree-decomposition of a graph $G$ is a rooted tree-decomposition $(\mathcal{T},\{X_{t}\}_{t\in V(\mathcal{T})})$ where $\mathcal{T}$ is rooted at $r$ with $X_{r}=\emptyset$, and each node of $\mathcal{T}$ belongs to one of the following types:
  • A leaf node $t$ is a leaf of $\mathcal{T}$ with $X_{t}=\emptyset$.
  • An introduce vertex node $t$ has one child $t_{1}$ such that $X_{t}=X_{t_{1}}\cup\{x\}$, for some $x\notin X_{t_{1}}$.
  • A forget node $t$ has one child $t_{1}$ such that $X_{t}=X_{t_{1}}\setminus\{x\}$, where $x\in X_{t_{1}}$.
  • An introduce edge node $t$ has one child $t^{\prime}$ with $X_{t}=X_{t^{\prime}}$, and is labeled with an edge $\{u,v\}\in E(G)$ such that $\{u,v\}\subseteq X_{t}$. We say that the edge $\{u,v\}$ is introduced at $t$;
  • A join node $t$ has two children $t_{1}$ and $t_{2}$ such that $X_{t}=X_{t_{1}}=X_{t_{2}}$.
Without loss of generality, we assume that a nice tree-decomposition of width tw is provided. If not, one can construct an optimum tree decomposition in time $2^{\mathcal{O}(\text{tw}^{2})}n^{\mathcal{O}(1)}$~[39].
Recall that we denote $G\cup H$ as the graph with vertex set $V(G)$ and edge set $E(G)\cup E(H)$. The following lemma enables us to work with tree decompositions of $G\cup H$ instead of just $G$, which is crucial for our formulation.
Lemma 3.1 ([41], Lemma 13).
If $H$ is a contraction of $G$, then $\textsf{tw}(G\cup H)\leq 2\textsf{tw}(G)$.
Henceforth, we fix a nice tree decomposition $\mathcal{T}$ of $G\cup H$. This choice is crucial because it ensures that for any edge $(u,v)\in E(H)$, there exists some bag $X_{t}$ such that $u,v\in X_{t}$, a condition that is not guaranteed in a tree decomposition of $G$ alone. For each node $t\in\mathcal{T}$, we define $G_{t}$ and $H_{t}$ to be the subgraphs of $G$ and $H$, respectively, induced by the union of all bags in the subtree of $\mathcal{T}$ rooted at $t$.
We now introduce the key definitions for the DP-state at a node $t$ of the tree-decomposition to convey the main ideas of the algorithm. The formal proof follows.
3.1.1 Witness Structures and Functional Representation
Assume that $H$ is a contraction of $G$, and let $\mathcal{W}=\{W_{1},\dots,W_{|V(H)|}\}$ denote a witness structure for this contraction. We represent $\mathcal{W}$ functionally using a pair of auxiliary functions: a parent function and an originator function.
Definition 3.3 (Parent function).
Given a witness structure $\mathcal{W}$, a function $\sigma^{\prime}:V(G)\to V(G)$ is a parent function with respect to $\mathcal{W}$ if there exists a collection of rooted trees $\{T_{i}\}_{i=1}^{k}$, where each $T_{i}$ is a spanning tree of $G[W_{i}]$ rooted at the unique vertex of $W_{i}\cap V(H)$, such that:
  1. 1.
    For all $u\in W_{i}\setminus V(H)$, $\sigma^{\prime}(u)$ is the parent of $u$ in $T_{i}$; and
  2. 2.
    For the unique vertex $u\in W_{i}\cap V(H)$, $\sigma^{\prime}(u)=u$.
Definition 3.4 (Originator function).
Given a witness structure $\mathcal{W}$, a function $\tau^{\prime}:V(G)\to V(H)$ is an originator function with respect to $\mathcal{W}$ if for every $u\in W_{i}$, $\tau^{\prime}(u)$ is the unique vertex in $W_{i}\cap V(H)$.
Note that any valid originator function $\tau^{\prime}$ defines a partition of $V(G)$ into the witness sets $\mathcal{W}$. Also, for any vertex $u\in V(G)$, there exists a bag that contains both $u$ and $\sigma^{\prime}(u)$, but not necessarily $\tau^{\prime}(u)$.
A pair $(\sigma^{\prime},\tau^{\prime})$ is said to be consistent if for every $u\in V(G)$ $\tau^{\prime}(u)=\tau^{\prime}(\sigma^{\prime}(u))$, i.e. $u$ and its parent $\sigma^{\prime}(u)$ belong to the same witness set. In the remainder of the discussion, we assume all parent-originator pairs to be consistent unless stated otherwise. Although $\tau^{\prime}$ can be derived from $\sigma^{\prime}$ by following parent pointers up to a vertex in $V(H)$, we maintain both functions explicitly for algorithmic clarity.
Definition 3.5 (Witness impression).
Every witness structure $\mathcal{W}$ corresponds to pair $(\sigma^{\prime},\tau^{\prime})$, called its witness-impression, where $\sigma^{\prime}$ is a parent function and $\tau^{\prime}$ is the corresponding originator function.
Definition 3.6 (Seed impression).
Let $(\sigma^{\prime},\tau^{\prime})$ be an impression of a witness structure. A pair of partial functions $(\sigma,\tau)$, with $\sigma:V(G_{t})\to V(G_{t})\cup\{\star\}$ and $\tau:V(G_{t})\to V(H_{t})\cup\{\star\}$, is a seed impression of $(\sigma^{\prime},\tau^{\prime})$ at $t$ if:
  • For every vertex $u\in V(G_{t})\setminus X_{t}$, $\sigma(u)=\sigma^{\prime}(u)$ and $\tau(u)=\tau^{\prime}(u)$.
  • For a vertex $u\in X_{t}$, if $\sigma^{\prime}(u)\in V(G_{t})$ then $\sigma(u)=\sigma^{\prime}(u)$ otherwise $\sigma(u)=\star$. The function $\tau$ is derived from $\sigma$ by following parent pointers up to either a vertex in $V(H)$ or $\star$.
Note that $(\sigma,\tau)$ is not merely the restriction of $(\sigma^{\prime},\tau^{\prime})$ to $V(G_{t})$, since a vertex $u$ such that $\sigma^{\prime}(u)\notin V(G_{t})$ might satisfy $\tau^{\prime}(u)\in V(G_{t})$, but we still assign $\tau(u)=\star$. Note that in Figure 1, $\tau^{\prime}(u_{1})\in G_{t}$ but since $\sigma^{\prime}(u_{1})\notin G_{t}$ we assign $\tau(u_{1})=\star$.
We now classify each vertex in the bag $X_{t}$ into one of three types, according to the seed impression $(\sigma,\tau)$:
Definition 3.7 (Types of vertices).
Let $t\in V(\mathcal{T})$ be a node with bag $X_{t}$, and let $(\sigma,\tau)$ be a seed impression at $t$. Then each vertex $u\in X_{t}$ is assigned one of the following types:
  • Requester: $\sigma(u)\in V(G_{t})$ and $\tau(u)\in V(H_{t})$ (i.e., both are defined); $u$ is ready to request adjacency requirements in $H$.
  • Dependent: $\sigma(u)\in V(G_{t})$ but $\tau(u)=\star$; $u$ has been contracted into its parent and depends on the parent node’s decision regarding the partition.
  • Explorer: $\sigma(u)=\star$ (implying $\tau(u)=\star$); $u$ is still exploring who its parent will be.
We classify the pair of requester vertices into the following three parts.
Definition 3.8 (Requester pair classification).
Let $u,v\in X_{t}$ be requester vertices with $\tau(u)=h_{u}$, $\tau(v)=h_{v}$. We say the pair $(u,v)$ is
  • invalid (inv) if $(h_{u},h_{v})\notin E(H)$ but $\tau^{-1}(h_{u}),\tau^{-1}(h_{v})$ are adjacent in $G_{t}$,
  • unsatisfied (unsat) if $(h_{u},h_{v})\in E(H)$ and $\tau^{-1}(h_{u}),\tau^{-1}(h_{v})$ are not adjacent in $G_{t}$, and
  • satisfied (sat), otherwise.
For a vertex $u\in V(G)$, we define the set $\textsf{des}(u)\subseteq V(G)$ of descendants of $u$ with respect to a function $\sigma^{\prime}:V(G)\to V(G)$ as follows:
$\textsf{des}(u)=\left\{v\in V(G)\;\middle|\;\text{there exists an integer }\ell\geq 0\text{ such that }\sigma^{\prime(\ell)}(v)=u\right\},$
where $\sigma^{\prime(0)}(v)=v$, and for $\ell\geq 1$, $\sigma^{\prime(\ell)}(v)=\sigma^{\prime}(\sigma^{\prime(\ell-1)}(v))$. That is, $\textsf{des}(u)$ consists of all vertices $v$ for which repeated applications of $\sigma^{\prime}$ eventually reach $u$, including $u$ itself.
Definition 3.9 (Pseudo-adjacency).
Let $u,v$ be two vertices in $X_{t}$. Then, we say $u$ and $v$ are pseudo-adjacent if sets $\textsf{des}(u)$ and $\textsf{des}(v)$ are adjacent in $G_{t}$.
3.1.2 Signatures and Dynamic Programming States
Suppose the graph $G$ admits an $H$-witness structure $\mathcal{W}$, and let $(\sigma^{\prime},\tau^{\prime})$ be the corresponding witness-impression as specified in Definition 3.5. Let $T$ be a tree decomposition of the graph $G\cup H$, and let $t$ be a node of $T$.
Definition 3.10 (Partial Solution).
We define a partial solution at node $t$ as a tuple $P=(\sigma,\tau,\mathcal{R},\mathcal{A}),$ with the following components:
  • $(\sigma,\tau)$ is a seed impression of the witness-impression $(\sigma^{\prime},\tau^{\prime})$, restricted to the bag $X_{t}$ at $t$, as specified in Definition 3.6,
  • $\mathcal{R}\subseteq R\times R\times\{\textsf{inv},\textsf{sat},\textsf{unsat}\}$ is a set of labeled pairs, where $R\subseteq X_{t}$ denotes the set of requester vertices contained in $X_{t}$ as per Definition 3.7. Each element $(u,v,\ell)\in\mathcal{R}$ represents a label $\ell\in\{\textsf{inv},\textsf{sat},\textsf{unsat}\}$ assigned to the pair $(u,v)$, as per Definition 3.8,
  • $\mathcal{A}\subseteq X_{t}\times X_{t}$ records unordered pairs of vertices in $X_{t}$ that are pseudo-adjacent to each other, as per Definition 3.9.
We now refer to Figure 1 to provide intuition for the above definitions. The figure shows the graph $H$ together with a part of the tree decomposition of $G\cup H$ (we do not explicitly draw $G$). The rounded outlines indicate the witness structure for the contraction, whenever it exists: each group contains exactly one node of $H$, representing the collection of vertices that contract to it. Note that $\sigma(u_{2})=u_{1}$, $\tau(u_{2})=\star$, and $\tau(u_{1})=\sigma(u_{1})=\star$. Since the edge $(u_{3},u_{4})$ is an edge we processed earlier in the DP, we may store $(u_{7},u_{5},\textsf{sat})\in\mathcal{R}$ for $X_{t}$. On the other hand, because the edge $(u_{6},u_{5})$ has not yet been introduced, the edge $(h_{2},h_{3})$ does not appear in the current structure, so we have $(u_{2},u_{5},\textsf{unsat})$. Furthermore, observe that $(u_{8},u_{5})$ and $(u_{7},u_{5})$ belong to $\mathcal{A}$, whereas $(u_{9},u_{5})\notin\mathcal{A}$. In this way, all relevant sets are defined for the dynamic programming.
$u_{8}$$u_{7}$$u_{1}$$h_{2}$$u_{2}$$u_{9}$$u_{5}$$u_{3}$$h_{1}$$u_{4}$$h_{3}$$u_{6}$ $X_{t}$$G_{t}$RestofGRest~of~G
(a) Illustration of $G_{t}$ and $X_{t}$
$h_{1}$$h_{2}$$h_{3}$
(b) Graph $H$
Figure 1: Depicting definitions relevant to the algorithm
We now provide a formal definition for the same.
Note that, as it is standard in the case of dynamic programming on tree decomposition, we will not compute all the partial solution at node $t$ but rather just store unique signatures of all the partial solutions at $t$. The signature $S$ of a partial solution $P=(\sigma,\tau,\mathcal{R},\mathcal{A})$ is $(\sigma|_{X_{t}},\tau|_{X_{t}},\mathcal{R},\mathcal{A})$.
States of the Dynamic Programming Algorithms
Definition 3.11 (Signature).
A state of dynamic programming table will be a tuple $S=(\sigma,\tau,\mathcal{R},\mathcal{A})$ such that
  • $\sigma:X_{t}\to X_{t}\cup\{\star\}$ (which will emulate a parent function),
  • $\tau:X_{t}\to V(H_{t})\cup\{\star\}$ (which will emulate an originator function),
  • $\mathcal{R}\subseteq R\times R\times\{\textsf{inv},\textsf{sat},\textsf{unsat}\}$ is a set of labeled pairs representing a label $\ell\in\{\textsf{inv},\textsf{sat},\textsf{unsat}\}$ assigned to the pair $(u,v)$, as per Definition 3.8, and
  • $\mathcal{A}$ is the collection of pair of vertices in $X_{t}$ that are pseudo-adjacent with each other.
The condition on functions ensures that at present $\sigma$ and $\tau$ are valid candidate to be a parent function and an originator function.
For a node $t$ in tree decomposition $T$, we construct a dynamic-programming algorithm that sets the value $\textbf{d}[S,t]$ to True if and only if $S$ is a signature of some partial solution $P$ at node $t$. Otherwise, we set it to False. Initially, all entries are set to False. For any signature $S=(\sigma,\tau,\mathcal{R},\mathcal{A})$, if $\mathcal{R}$ contains an entry of the type $(u,v,\textsf{inv})$ then $\textbf{d}[S,t]$ is set to False and it will never be changed.
3.1.3 Updating Dynamic Programming Tables
We now proceed to describe a recursive algorithm used to update entries in the dynamic programming table for each node type in a nice tree-decomposition.
Leaf Node
Let $t$ be a leaf node. Then $X_{t}=\emptyset$ and the valid partial solution $P$ is empty functions $\sigma,\tau,\mathcal{R}$ since their domain is a subset of $X_{t}$ which is trivially empty. Hence, the only valid signature consists of four empty sets.
Lemma 3.2.
If $t$ is a leaf node then set $\textbf{d}[S,t]=\texttt{True}$ for $S=(\emptyset,\emptyset,\emptyset,\emptyset)$, and $\textbf{d}[S^{\prime},t]=\texttt{False}$ for every other tuple $S^{\prime}$.
Introduce Vertex Node
Let $t$ be an introduce vertex node in the tree decomposition $T$, with child $t^{\prime}$, such that $X_{t}=X_{t^{\prime}}\cup\{x\}$ for some vertex $x\in V(G)\setminus V(G_{t^{\prime}})$. Let $S=(\sigma,\tau,\mathcal{R},\mathcal{A})$ be a signature at node $t$, and let $S^{\prime}=(\sigma^{\prime},\tau^{\prime},\mathcal{R}^{\prime},\mathcal{A}^{\prime})$ be a signature at node $t^{\prime}$.
Definition 3.12.
We say that $S$ and $S^{\prime}$ are introduce-vertex compatible if the following conditions hold:
  • ($C1$)
    For every $y\in X_{t}\setminus\{x\}$, we have $\sigma(y)=\sigma^{\prime}(y)$ and $\tau(y)=\tau^{\prime}(y)$,
  • ($C2$)
    $\mathcal{R}^{\prime}\subseteq\mathcal{R}$ and $\mathcal{A}^{\prime}=\mathcal{A}$.
Let $I(S)$ denote the set of all signatures at $t^{\prime}$ that are introduce-vertex compatible with $S$.
Lemma 3.3.
Let $t$ be an introduce vertex node with child $t^{\prime}$, and let $X_{t}=X_{t^{\prime}}\cup\{x\}$ and $x\notin V(H)$. Let $S=(\sigma,\tau,\mathcal{R},\mathcal{A})$ be a signature at node $t$. Assign $\mathbf{d}[S,t]=\texttt{True}$ if and only if the following conditions are satisfied:
  1. (i)
    $\sigma(x)=\star$ and $\tau(x)=\star$,
  2. (ii)
    there exists a signature $S^{\prime}=(\sigma^{\prime},\tau^{\prime},\mathcal{R}^{\prime},\mathcal{A}^{\prime})\in I(S)$ such that $\mathbf{d}[S^{\prime},t^{\prime}]=\texttt{True}$,
  3. (iii)
    $\mathcal{R}=\mathcal{R}^{\prime}$ and $\mathcal{A}=\mathcal{A}^{\prime}$, i.e., no new elements are introduced in either of the sets.
Proof.
$(\Leftarrow)$ Assume the three conditions of the lemma are met. By condition (ii), there exists a valid signature $S^{\prime}=(\sigma^{\prime},\tau^{\prime},\mathcal{R}^{\prime},\mathcal{A}^{\prime})$ at the child node $t^{\prime}$, which means there is a valid partial solution $P^{\prime}=(\sigma^{\prime}_{p},\tau^{\prime}_{p},\mathcal{R}^{\prime}_{p},\mathcal{A}^{\prime}_{p})$ for the subgraph $G_{t^{\prime}}$ that corresponds to $S^{\prime}$. We construct a new partial solution $P=(\sigma_{p},\tau_{p},\mathcal{R}_{p},\mathcal{A}_{p})$ for the subgraph $G_{t}$ by extending $P^{\prime}$ to include vertex $x$. For all $u\in V(G_{t^{\prime}})$, we set $\sigma_{p}(u)=\sigma^{\prime}_{p}(u)$ and $\tau_{p}(u)=\tau^{\prime}_{p}(u)$. For the new vertex $x\notin V(H)$, its parent must lie outside the current subtree. Thus, we set $\sigma_{p}(x)=\star$ and $\tau_{p}(x)=\star$, consistent with condition (i). Since no new edges are introduced involving $x$, we set $\mathcal{R}_{p}=\mathcal{R}^{\prime}_{p}$ and $\mathcal{A}_{p}=\mathcal{A}^{\prime}_{p}$, consistent with condition (iii). This constructed tuple $P$ is a valid partial solution for $G_{t}$, and its signature is precisely $S$. Therefore, $\mathbf{d}[S,t]$ is set to True.
$(\Rightarrow)$ Now, assume that $\mathbf{d}[S,t]=\text{true}$. This means there exists a valid partial solution $P=(\sigma_{p},\tau_{p},\mathcal{R}_{p},\mathcal{A}_{p})$ for $G_{t}$ whose signature is $S$. We can define a partial solution $P^{\prime}$ for the smaller subgraph $G_{t^{\prime}}$ by simply restricting $P$. For all $u\in V(G_{t^{\prime}})$, we set $\sigma^{\prime}_{p}(u)=\sigma_{p}(u)$, $\tau^{\prime}_{p}(u)=\tau_{p}(u)$, $\mathcal{R}^{\prime}_{p}=\mathcal{R}_{p}$, and $\mathcal{A}^{\prime}_{p}=\mathcal{A}_{p}$. This new tuple $P^{\prime}$ is a valid partial solution for $G_{t^{\prime}}$, and its signature $S^{\prime}$ must exist in the DP table, meaning $\mathbf{d}[S^{\prime},t^{\prime}]=\text{true}$. This proves condition (ii). Since $x\notin V(H)$ and is a new, isolated vertex, it cannot be a parent to any other vertex in the subtree. This requires that in any valid partial solution, $\sigma_{p}(x)=\star$ and $\tau_{p}(x)=\star$, which proves condition (i). Finally, because no new edges or relationships were formed, the sets of requester pairs and pseudo-adjacencies remain unchanged from $t^{\prime}$ to $t$. Thus, $\mathcal{R}=\mathcal{R}^{\prime}$ and $\mathcal{A}=\mathcal{A}^{\prime}$, which proves condition (iii). ∎
Lemma 3.4.
Let $t$ be an introduce vertex node with child $t^{\prime}$, and let $X_{t}=X_{t^{\prime}}\cup\{x\}$ and $x\in V(H)$. Let $S=(\sigma,\tau,\mathcal{R},\mathcal{A})$ be a signature at node $t$. Assign the value of $\mathbf{d}[S,t]=\texttt{True}$ if and only if the following conditions are met:
  1. (i)
    $\sigma(x)=x$ and $\tau(x)=x$,
  2. (ii)
    there exists a signature $S^{\prime}=(\sigma^{\prime},\tau^{\prime},\mathcal{R}^{\prime},\mathcal{A}^{\prime})\in I(S)$ such that $\mathbf{d}[S^{\prime},t^{\prime}]=\texttt{True}$,
  3. (iii)
    $\mathcal{A}=\mathcal{A}^{\prime}$, and every element in $\mathcal{R}\setminus\mathcal{R}^{\prime}$ is of the form
    • $(x,u,\textsf{unsat})$ for some $u\in X_{t}$ such that $x$ and $\tau(u)$ are adjacent in $H$, or
    • $(x,u,\textsf{sat})$ for some $u\in X_{t}$ such that $x$ and $\tau(u)$ are not adjacent in $H$.
    The above condition implies that the only tuple introduced in $\mathcal{R}$ contains $x$ in it.
Proof.
$(\Leftarrow)$ Assume the three conditions of the lemma are met. By condition (ii), there exists a valid signature $S^{\prime}=(\sigma^{\prime},\tau^{\prime},\mathcal{R}^{\prime},\mathcal{A}^{\prime})$ at the child node $t^{\prime}$, which means there is a valid partial solution $P^{\prime}=(\sigma^{\prime}_{p},\tau^{\prime}_{p},\mathcal{R}^{\prime}_{p},\mathcal{A}^{\prime}_{p})$ for the subgraph $G_{t^{\prime}}$. We construct a new partial solution $P=(\sigma_{p},\tau_{p},\mathcal{R}_{p},\mathcal{A}_{p})$ for the subgraph $G_{t}$ by extending $P^{\prime}$ to include vertex $x$. For all $u\in V(G_{t^{\prime}})$, we set $\sigma_{p}(u)=\sigma^{\prime}_{p}(u)$ and $\tau_{p}(u)=\tau^{\prime}_{p}(u)$. For the new vertex $x\in V(H)$, it must be the root of its own witness set. Thus, we set $\sigma_{p}(x)=x$ and $\tau_{p}(x)=x$, which is consistent with condition (i). The pseudo-adjacency set is unchanged as no new edges from $G$ are introduced, so we set $\mathcal{A}_{p}=\mathcal{A}^{\prime}_{p}$, consistent with condition (iii). As $x$ is now a requester vertex, its relationships with other requesters in the bag must be recorded. Based on condition (iii), we define $\mathcal{R}_{p}=\mathcal{R}^{\prime}_{p}\cup\{(x,u,\textsf{unsat})\mid(x,\tau(u))\in E(H)\}\cup\{(x,u,\textsf{sat})\mid(x,\tau(u))\notin E(H)\}$. The constructed tuple $P$ is a valid partial solution for $G_{t}$, and its signature is precisely $S$. Therefore, $\mathbf{d}[S,t]$ is set to True.
$(\Rightarrow)$ Now, assume that $\mathbf{d}[S,t]=\text{true}$. This means there exists a valid partial solution $P=(\sigma_{p},\tau_{p},\mathcal{R}_{p},\mathcal{A}_{p})$ for $G_{t}$ whose signature is $S$. We can define a partial solution $P^{\prime}$ for the smaller subgraph $G_{t^{\prime}}$ by simply restricting $P$. For all $u\in V(G_{t^{\prime}})$, we set $\sigma^{\prime}_{p}(u)=\sigma_{p}(u)$ and $\tau^{\prime}_{p}(u)=\tau_{p}(u)$. The sets $\mathcal{R}^{\prime}_{p}$ and $\mathcal{A}^{\prime}_{p}$ are simply the projections of $\mathcal{R}_{p}$ and $\mathcal{A}_{p}$ onto pairs of vertices in $X_{t^{\prime}}$. The resulting tuple $P^{\prime}$ is a valid partial solution for $G_{t^{\prime}}$ since it omits the new vertex $x$. Its signature $S^{\prime}$ must exist in the DP table, meaning $\mathbf{d}[S^{\prime},t^{\prime}]=\text{true}$. This proves condition (ii). Since $x\in V(H)$, by the properties of a valid witness structure, $x$ must be the root of its witness tree and its own originator. This means $\sigma(x)=x$ and $\tau(x)=x$, which proves condition (i). Finally, as no edges from $G$ are introduced, the pseudo-adjacency set for pairs in $X_{t}$ remains unchanged from $t^{\prime}$ to $t$. Any new entries in the requester set $\mathcal{R}$ must involve the new vertex $x$ and must be correctly classified based on its relationships to other requesters, as captured in condition (iii). ∎
Forget Vertex Node.
Let $t$ be a forget vertex node in the tree decomposition $T$, with child $t^{\prime}$, such that $X_{t}=X_{t^{\prime}}\setminus\{x\}$ for some vertex $x\in X_{t^{\prime}}$. Let $S=(\sigma,\tau,\mathcal{R},\mathcal{A})$ be a signature at node $t$, and let $S^{\prime}=(\sigma^{\prime},\tau^{\prime},\mathcal{R}^{\prime},\mathcal{A}^{\prime})$ be a signature at node $t^{\prime}$.
Definition 3.13.
We say that $S$ and $S^{\prime}$ are forget-vertex compatible if the following conditions hold:
  • (C1) The vertex $x$ is not an explorer with respect to $\sigma^{\prime}$, as per Definition 3.7.
  • (C2) For every $y\in X_{t^{\prime}}\setminus\{x\}$, we have $\sigma(y)=\sigma^{\prime}(y)$ and $\tau(y)=\tau^{\prime}(y)$.
  • (C3) $\mathcal{R}$ is the restriction of $\mathcal{R}^{\prime}$ to pairs not involving $x$, and $\mathcal{A}$ is the restriction of $\mathcal{A}^{\prime}$ to pairs not involving $x$.
Let $F(S)$ denote the set of all signatures at $t^{\prime}$ that are forget-vertex compatible with $S$.
Lemma 3.5.
Let $t$ be a forget vertex node with child $t^{\prime}$, and let $X_{t}=X_{t^{\prime}}\setminus\{x\}$. Let $S=(\sigma,\tau,\mathcal{R},\mathcal{A})$ be a signature at node $t$. Then, $\mathbf{d}[S,t]=\texttt{True}$ if and only if the following conditions hold:
  1. (i)
    There exists a signature $S^{\prime}=(\sigma^{\prime},\tau^{\prime},\mathcal{R}^{\prime},\mathcal{A}^{\prime})\in F(S)$ such that $\mathbf{d}[S^{\prime},t^{\prime}]=\texttt{True}$.
  2. (ii)
    The requester set $\mathcal{R}^{\prime}$ contains no entry of the form $(x,u,\textsf{unsat})$ for any $u\in X_{t^{\prime}}\setminus\{x\}$.
Proof.
$(\Leftarrow)$ Assume the two conditions of the lemma are met. By condition (i), there exists a valid signature $S^{\prime}$ at $t^{\prime}$ corresponding to a valid partial solution $P^{\prime}$ for the subgraph $G_{t^{\prime}}$. We construct a new partial solution $P$ for $G_{t}$ by restricting $P^{\prime}$ to $V(G)\setminus\{x\}$. Since $P^{\prime}$ is valid, all adjacency requirements imposed on $G_{t^{\prime}}$ must hold. Condition (ii) explicitly guarantees that the vertex $x$ does not have any unsat adjacency requirements with any other vertex $u\in X_{t}$. Because $x$ is not an explorer (C1), all required parent/originator relations for $x$ must be contained within $G_{t^{\prime}}$. By C3, the signature $S$ is precisely the valid signature $S^{\prime}$ with all information about $x$ correctly removed. Since all constraints involving $x$ are resolved within $G_{t^{\prime}}$ and $P^{\prime}$ is valid, the restriction $P$ is a valid partial solution for $G_{t}$. Therefore, $\mathbf{d}[S,t]$ is set to True.
$(\Rightarrow)$ Now, assume that $\mathbf{d}[S,t]=\text{true}$. This means there exists a valid partial solution $P$ for $G_{t}$ whose signature is $S$. This solution $P$ must be the restriction of some valid partial solution $P^{\prime}$ for $G_{t^{\prime}}$. This proves the existence of a compatible signature $S^{\prime}$ such that $\mathbf{d}[S^{\prime},t^{\prime}]=\text{true}$, proving condition (i). For $P^{\prime}$ to be a valid partial solution, all requirements within $G_{t^{\prime}}$ must be satisfied. If $\mathcal{R}^{\prime}$ contained an entry $(x,u,\textsf{unsat})$, the required edge in $H$ between $\tau(x)$ and $\tau(u)$ would not have been realized by the edges in $G_{t^{\prime}}$, contradicting the validity of $P^{\prime}$. Thus, no unsat entry involving $x$ can exist in $\mathcal{R}^{\prime}$. This proves condition (ii). ∎
Introduce Edge Node
Let $t$ be an introduce edge node with child $t^{\prime}$, so that $X_{t}=X_{t^{\prime}}$, and let the edge $(u,v)\in E(G)$ be introduced at $t$, with $u,v\in X_{t}$. Let $S=(\sigma,\tau,\mathcal{R},\mathcal{A})$ be a candidate signature at $t$.
Definition 3.14.
We say that $S$ is introduce-edge compatible with respect to its counterpart at $t^{\prime}$, denoted $S^{\prime}=(\sigma^{\prime},\tau^{\prime},\mathcal{R}^{\prime},\mathcal{A}^{\prime})$, if the following hold:
  • (C1) For every $x\in X_{t}\setminus\{u,v\}$, we have $\sigma(x)=\sigma^{\prime}(x)$ and $\tau(x)=\tau^{\prime}(x)$.
  • (C2) At most one of the following statements is true for the values of $\sigma$ and $\tau$ at $u$ and $v$:
    • (C2.1) No change in values. Formally, $\sigma^{\prime}(u)=\sigma(u)$, $\tau^{\prime}(u)=\tau(u)$ and $\sigma^{\prime}(v)=\sigma(v)$, $\tau^{\prime}(v)=\tau(v)$.
    • (C2.2) $v$’s set merges into $u$’s set. Formally, $\sigma(u)=\sigma^{\prime}(u)$, $\tau(u)=\tau^{\prime}(u)$, $\sigma(v)\neq\sigma^{\prime}(v)$, and $\tau(v)=\tau(u)$. This change represents $u$ becoming the parent of $v$ or the root of $v$’s witness tree.
    • (C2.3) $u$’s set merges into $v$’s set. Formally, $\sigma(v)=\sigma^{\prime}(v)$, $\tau(v)=\tau^{\prime}(v)$, $\sigma(u)\neq\sigma^{\prime}(u)$, and $\tau(u)=\tau(v)$.
  • (C3) $\mathcal{R}^{\prime}\setminus\mathcal{R}\subseteq\{(u,v,\textsf{unsat}),(v,u,\textsf{unsat})\}$ and $\mathcal{A}^{\prime}\subseteq\mathcal{A}$.
Let $I_{e}(S)$ denote the set of all signatures at $t^{\prime}$ that are introduce-edge compatible with $S$.
Lemma 3.6.
Let $t$ be an introduce edge node where edge $(u,v)$ is introduced, with $u,v\in V(H)$. A signature $S$ is valid at node $t$ if and only if there exists a signature $S^{\prime}\in I_{e}(S)$ such that $\mathbf{d}[S^{\prime},t^{\prime}]=\texttt{True}$, and the following conditions hold:
  1. (i)
    Condition C2.1 holds.
  2. (ii)
    The pseudo-adjacency set is unchanged: $\mathcal{A}=\mathcal{A}^{\prime}$.
  3. (iii)
    The requester set is updated only for $(u,v)$: $\mathcal{R}=\mathcal{R}^{\prime}\setminus\{(u,v,\textsf{unsat})\}\cup\{(u,v,\textsf{sat})\}$, if $(u,v,\textsf{unsat})\in\mathcal{R}^{\prime}$. Otherwise $\mathcal{R}=\mathcal{R}^{\prime}$.
Proof.
($\Leftarrow$) Assume the conditions are met. Condition (i) must hold since $u,v\in V(H)$, meaning they must be the roots of their witness sets, which cannot merge due to a $G$-edge. By the premise, a valid partial solution $P^{\prime}$ exists for $G_{t^{\prime}}$. We construct $P$ for $G_{t}$ by adding $(u,v)$. Since the functions $\sigma_{p},\tau_{p}$ and $\mathcal{A}_{p}$ are unchanged (condition (i) and (ii)), the only effect is to satisfy the $H$-requirement between $\tau(u)$ and $\tau(v)$. Condition (iii) correctly formalizes this satisfaction, updating an unsat requirement to sat in $\mathcal{R}$. Thus $P$ is a valid partial solution for $G_{t}$.
($\Rightarrow$) If $\mathbf{d}[S,t]=\text{true}$, a valid $P$ exists, derived from a valid $P^{\prime}$ at $t^{\prime}$. Since $u,v\in V(H)$, the edge $(u,v)$ cannot cause a witness set merge, so the functions must remain the same, proving condition (i). The edge is a $G$-edge, not a pseudo-adjacency, so $\mathcal{A}$ remains $\mathcal{A}^{\prime}$, proving condition (ii). Its only effect is on the $H$-contraction requirement, which means if $(u,v,\textsf{unsat})$ was in $\mathcal{R}^{\prime}$, it must be resolved to sat in $\mathcal{R}$. If it was already sat or inv, $\mathcal{R}$ must be the same as $\mathcal{R}^{\prime}$. This verifies condition (iii). ∎
Lemma 3.7.
Let $t$ be an introduce edge node where edge $(u,v)$ is introduced, with $u\in V(H)$ and $v\notin V(H)$. A signature $S$ is valid at node $t$ if and only if there exists a signature $S^{\prime}\in I_{e}(S)$ such that $\mathbf{d}[S^{\prime},t^{\prime}]=\texttt{True}$, and one of the following holds:
  1. (i)
    Condition C2.1 holds and $\mathcal{R}=\mathcal{R}^{\prime}$ and $\mathcal{A}=\mathcal{A}^{\prime}\cup\{(u,v)\}$.
  2. (ii)
    Condition C2.2 holds and $\mathcal{R}=\mathcal{R}^{\prime}$ and $\mathcal{A}=\mathcal{A}^{\prime}$.
Proof.
($\Leftarrow$) Assume one of the conditions holds. A valid $S^{\prime}$ exists by the premise.
  1. 1.
    If (i) holds, the functions are preserved (C2.1), meaning no witness set merge occurs. The edge $(u,v)$ is thus recorded as a new pseudo-adjacency in $\mathcal{A}$, forming a valid $P$ for $G_{t}$.
  2. 2.
    If (ii) holds, the functions change according to C2.2, meaning the witness set containing $v$ merges into $u$’s set. The edge $(u,v)$ is absorbed into the new witness structure and does not become a pseudo-adjacency, so $\mathcal{A}$ remains $\mathcal{A}^{\prime}$. This forms a valid $P$ for $G_{t}$.
($\Rightarrow$) If $\mathbf{d}[S,t]=\text{true}$, a valid $P$ exists. Since $u\in V(H)$ and $v\notin V(H)$, $u$ must be a requester and $v$ cannot be the root of a witness set. The new edge $(u,v)$ must either form a pseudo-adjacency (case (i)) or cause $v$’s set to be absorbed by $u$’s (case (ii)). Case (i) requires function stability (C2.1). Case (ii) requires the witness set merge, which corresponds to the $\sigma$ and $\tau$ changes in C2.2. These two cases are exhaustive and mutually exclusive for a valid solution. ∎
Lemma 3.8.
Let $t$ be an introduce edge node where edge $(u,v)$ is introduced, with $u\notin V(H)$ and $v\in V(H)$. A signature $S$ is valid at node $t$ if and only if there exists a signature $S^{\prime}\in I_{e}(S)$ such that $\mathbf{d}[S^{\prime},t^{\prime}]=\texttt{True}$, and one of the following holds:
  1. (i)
    Condition C2.1 holds and $\mathcal{A}=\mathcal{A}^{\prime}\cup\{(u,v)\}$.
  2. (ii)
    Condition C2.3 holds and $\mathcal{R}=\mathcal{R}^{\prime}$ and $\mathcal{A}=\mathcal{A}^{\prime}$.
Proof.
This is symmetric to Lemma 3.7, with $u$ and $v$ roles swapped. Case (i) is for pseudo-adjacency (no merge), and Case (ii) is for $u$’s set merging into $v$’s set (C2.3). The proof follows the same logic. ∎
Lemma 3.9.
Let $t$ be an introduce edge node where edge $(u,v)$ is introduced, with $u\notin V(H)$ and $v\notin V(H)$. A signature $S$ is valid at node $t$ if and only if there exists a signature $S^{\prime}\in I_{e}(S)$ such that $\mathbf{d}[S^{\prime},t^{\prime}]=\texttt{True}$, and one of the following holds:
  1. (i)
    Condition C2.1 holds and $\mathcal{A}=\mathcal{A}^{\prime}\cup\{(u,v)\}$.
  2. (ii)
    Condition C2.2 holds and $\mathcal{R}=\mathcal{R}^{\prime}$ and $\mathcal{A}=\mathcal{A}^{\prime}$.
  3. (iii)
    Condition C2.3 holds and $\mathcal{R}=\mathcal{R}^{\prime}$ and $\mathcal{A}=\mathcal{A}^{\prime}$.
Proof.
This case combines the possibilities from the previous two lemmas. Since neither $u$ nor $v$ are $H$-vertices, the edge $(u,v)$ can either: (i) be a pseudo-adjacency (functions unchanged, C2.1), or (ii) cause $v$’s set to merge into $u$’s (C2.2), or (iii) cause $u$’s set to merge into $v$’s (C2.3). In the merge cases ((ii) and (iii)), the edge is absorbed into the witness structure, leaving $\mathcal{A}$ unchanged. The proof follows the same construction and existence logic as the previous cases. ∎
Join Node
Let $t$ be a join node in the tree decomposition $T$, with children $t_{1}$ and $t_{2}$, such that $X_{t}=X_{t_{1}}=X_{t_{2}}$. Let $S=(\sigma,\tau,\mathcal{R},\mathcal{A})$ be a signature at node $t$, and let $S_{1}=(\sigma_{1},\tau_{1},\mathcal{R}_{1},\mathcal{A}_{1})$ and $S_{2}=(\sigma_{2},\tau_{2},\mathcal{R}_{2},\mathcal{A}_{2})$ be signatures at nodes $t_{1}$ and $t_{2}$ respectively.
Definition 3.15.
We say that $S$, $S_{1}$, and $S_{2}$ are join-node compatible if the following conditions hold:
  • (C1) For every $y\in X_{t}$, we have $\sigma(y)=\sigma_{1}(y)=\sigma_{2}(y)$ and $\tau(y)=\tau_{1}(y)=\tau_{2}(y)$.
  • (C2) The signature sets are combined by union: $\mathcal{R}=\mathcal{R}_{1}\cup\mathcal{R}_{2}$ and $\mathcal{A}=\mathcal{A}_{1}\cup\mathcal{A}_{2}$.
Let $J(S)$ denote the set of all pairs of signatures $(S_{1},S_{2})$ at $t_{1}$ and $t_{2}$ respectively, that satisfy (C1) and (C2) with $S$.
Lemma 3.10.
Let $t$ be a join node with children $t_{1}$ and $t_{2}$, and let $X_{t}=X_{t_{1}}=X_{t_{2}}$. Let $S=(\sigma,\tau,\mathcal{R},\mathcal{A})$ be a signature at node $t$. Then, $\mathbf{d}[S,t]=\texttt{True}$ if and only if there exists a pair of signatures $(S_{1},S_{2})\in J(S)$, where $S_{1}=(\sigma_{1},\tau_{1},\mathcal{R}_{1},\mathcal{A}_{1})$ is a signature at $t_{1}$ and $S_{2}=(\sigma_{2},\tau_{2},\mathcal{R}_{2},\mathcal{A}_{2})$ is a signature at $t_{2}$, such that:
  1. (i)
    $\mathbf{d}[S_{1},t_{1}]=\texttt{True}$ and $\mathbf{d}[S_{2},t_{2}]=\texttt{True}$.
  2. (ii)
    The sets $\mathcal{A}$ and $\mathcal{R}$ are formed by combining the corresponding sets from $S_{1}$ and $S_{2}$ according to the following rules for every pair $(u,v)\in X_{t}\times X_{t}$:
    • $\mathcal{A}=\mathcal{A}_{1}\cup\mathcal{A}_{2}$.
    • $\mathcal{R}$ is the union of all combined requester entries, $\mathcal{R}=\bigcup_{u,v\in X_{t}}\mathcal{R}(u,v)$, where $\mathcal{R}(u,v)$ is determined by the $\oplus$ and $\otimes$ operators:
      • Requester Merge ($\oplus$): For $(u,v,\ell_{1})\in\mathcal{R}_{1}$ and $(u,v,\ell_{2})\in\mathcal{R}_{2}$:
[…]
      • Pseudo-Adjacency Satisfies ($\otimes$): For $(u,v,\ell_{1})\in\mathcal{R}_{1}$ and $(u,v)\in\mathcal{A}_{2}$:
        $\ell_{1}\otimes\mathcal{A}_{2}=\begin{cases}\textsf{sat}&\text{if }\ell_{1}=\textsf{unsat}\\ \ell_{1}&\text{if }\ell_{1}\in\{\textsf{sat},\textsf{inv}\}\end{cases}$
        (A symmetric rule applies for $\mathcal{R}_{2}\otimes\mathcal{A}_{1}$.)
      The resulting entry for $(u,v)$ in $\mathcal{R}$ is $\ell_{1}\oplus\ell_{2}$ (if both exist), or the result of the $\otimes$ operations, prioritizing sat over unsat and unsat over non-existence, and rejecting inv.
Proof.
$(\Leftarrow)$ Assume conditions (i) and (ii) hold. By condition (i), there exist valid partial solutions $P_{1}=(\sigma_{p1},\tau_{p1},\mathcal{R}_{p1},\mathcal{A}_{p1})$ for $G_{t_{1}}$ and $P_{2}=(\sigma_{p2},\tau_{p2},\mathcal{R}_{p2},\mathcal{A}_{p2})$ for $G_{t_{2}}$. We construct a new partial solution $P=(\sigma_{p},\tau_{p},\mathcal{R}_{p},\mathcal{A}_{p})$ for $G_{t}=G_{t_{1}}\cup G_{t_{2}}$. The functions $\sigma_{p}$ and $\tau_{p}$ are defined piece-wise by the union of $\sigma_{p1},\tau_{p1}$ and $\sigma_{p2},\tau_{p2}$. Since $X_{t}=X_{t_{1}}=X_{t_{2}}$, condition (C1) ensures the functions agree on $X_{t}$, guaranteeing that the union is a valid structure for the witness sets. The set of edges $E(G_{t})=E(G_{t_{1}})\cup E(G_{t_{2}})$.
  • Pseudo-Adjacency ($\mathcal{A}$): The pseudo-adjacencies in $P$ are exactly the union of those in $P_{1}$ and $P_{2}$, as $G_{t_{1}}$ and $G_{t_{2}}$ share no edges. Condition (ii) enforces $\mathcal{A}=\mathcal{A}_{1}\cup\mathcal{A}_{2}$.
  • Requester Set ($\mathcal{R}$): The requester set tracks $H$-requirements between the originators of vertices in $X_{t}$. A requirement is satisfied in $G_{t}$ if it is satisfied in $G_{t_{1}}$ or $G_{t_{2}}$, or if the necessary edge is provided by a pseudo-adjacency in the other subgraph.
    • The $\oplus$ operator correctly determines the status based on the union of $G_{t_{1}}$ and $G_{t_{2}}$: an unsat requirement becomes sat if the edge is found in the other subgraph (represented by an sat in the other $\mathcal{R}$ set).
    • The $\otimes$ operator correctly models the satisfaction of an unsat requirement in one subgraph ($\mathcal{R}_{1}$ or $\mathcal{R}_{2}$) by a pseudo-adjacency recorded in the other subgraph ($\mathcal{A}_{2}$ or $\mathcal{A}_{1}$).
    Condition (ii) correctly computes the final status for $\mathcal{R}$ ensuring all constraints are properly inherited and resolved.
Since $P_{1}$ and $P_{2}$ are valid, and $P$ is constructed consistently with the union rules, $P$ is a valid partial solution for $G_{t}$. The signature of $P$ is $S$, thus $\mathbf{d}[S,t]=\texttt{True}$.
$(\Rightarrow)$ Assume $\mathbf{d}[S,t]=\text{true}$. There exists a valid partial solution $P$ for $G_{t}$ whose signature is $S$. By the definition of $G_{t}$ and tree decomposition, $P$ must be the result of combining two valid partial solutions $P_{1}$ for $G_{t_{1}}$ and $P_{2}$ for $G_{t_{2}}$.
  • Let $S_{1}$ and $S_{2}$ be the signatures of $P_{1}$ and $P_{2}$, respectively. Since $P_{1}$ and $P_{2}$ are valid, condition (i) must hold.
  • Since $P$ is a valid union of $P_{1}$ and $P_{2}$ and $X_{t}$ is the intersection, the function values must agree on $X_{t}$, which verifies (C1) and ensures $(S_{1},S_{2})\in J(S)$.
  • The pseudo-adjacency set $\mathcal{A}$ in $S$ must be the union of the pseudo-adjacencies recorded in $S_{1}$ and $S_{2}$ as the edges are disjoint, verifying $\mathcal{A}=\mathcal{A}_{1}\cup\mathcal{A}_{2}$.
  • The final status of any $H$-requirement $(u,v)$ in $\mathcal{R}$ must be the result of combining the status from $S_{1}$ and $S_{2}$, accounting for satisfaction via: (1) internal edges in $G_{t_{1}}$, (2) internal edges in $G_{t_{2}}$, or (3) pseudo-adjacencies in the opposite subgraph. This requires the exact combination defined by the $\oplus$ and $\otimes$ operators in condition (ii) to hold.
Since a valid $P$ exists, a pair $(S_{1},S_{2})$ satisfying all specified conditions must exist. ∎
3.1.4 Putting it all together.
We are now in a position to present the proof of Theorem 1 which states that Labeled Contraction admits an algorithm with running time $2^{\mathcal{O}(\textsf{tw}^{2})}\cdot|V(G)|^{\mathcal{O}(1)}$.
Proof.
The proof relies on the Dynamic Programming (DP) approach defined by the signatures $S=(\sigma,\tau,\mathcal{R},\mathcal{A})$ and the update laws (Lemmas 3.2 through 3.10). The overall running time is determined by the number of nodes in the tree decomposition, the size of the DP state space, and the time required for each transition.
At any node $t$ of the tree decomposition, the size of the bag $|X_{t}|$ is at most $\textsf{tw}+1$. The number of distinct signatures $S=(\sigma,\tau,\mathcal{R},\mathcal{A})$ stored at node $t$ is bounded by considering each component: The Parent Function $\sigma:X_{t}\to X_{t}\cup\{\star\}$ and the Originator Function $\tau:X_{t}\to V(H)\cup\{\star\}$ each contribute $2^{\mathcal{O}(\textsf{tw}\log\textsf{tw})}$ states. For $\tau$, this is because only $\mathcal{O}(\textsf{tw})$ vertices in $V(H)$ are relevant to the bag $X_{t}$. The Requester Set $\mathcal{R}$ encodes a status from $\{\textsf{sat},\textsf{unsat},\textsf{inv}\}$ for $\mathcal{O}(\textsf{tw}^{2})$ pairs, resulting in $3^{\mathcal{O}(\textsf{tw}^{2})}=2^{\mathcal{O}(\textsf{tw}^{2})}$ choices. The Pseudo-Adjacency Set $\mathcal{A}$ is binary for $\mathcal{O}(\textsf{tw}^{2})$ pairs, contributing $2^{\mathcal{O}(\textsf{tw}^{2})}$ choices. Multiplying these components yields the total number of distinct signatures:
$\text{Total States}=2^{\mathcal{O}(\textsf{tw}\log\textsf{tw})}\cdot 2^{\mathcal{O}(\textsf{tw}\log\textsf{tw})}\cdot 2^{\mathcal{O}(\textsf{tw}^{2})}\cdot 2^{\mathcal{O}(\textsf{tw}^{2})}=2^{\mathcal{O}(\textsf{tw}^{2})}$
The DP algorithm processes each of the $\mathcal{O}(|V(G)|)$ nodes in the tree decomposition. At Leaf, Introduce, and Forget Nodes, the transition involves iterating over the $2^{\mathcal{O}(\textsf{tw}^{2})}$ signatures of the single child node, with the compatibility checks and updates for each taking time polynomial in the bag size, $\mathcal{O}(\textsf{tw}^{\mathcal{O}(1)})$. At Join Nodes, the transition requires checking all pairs of compatible signatures $(S_{1},S_{2})$ from the two child nodes. Since there are $2^{\mathcal{O}(\textsf{tw}^{2})}$ signatures at each child, there are at most $\left(2^{\mathcal{O}(\textsf{tw}^{2})}\right)^{2}=2^{\mathcal{O}(\textsf{tw}^{2})}$ pairs to examine, and the combining operations ($\oplus$ and $\otimes$) for each pair take time $\mathcal{O}(\textsf{tw}^{\mathcal{O}(1)})$. Consequently, the time required to process any single node is bounded by $2^{\mathcal{O}(\textsf{tw}^{2})}\cdot\textsf{tw}^{\mathcal{O}(1)}=2^{\mathcal{O}(\textsf{tw}^{2})}$.
The overall running time of the algorithm is the number of nodes in the decomposition multiplied by the time per node:
$\text{Time}=\mathcal{O}(|V(G)|)\cdot 2^{\mathcal{O}(\textsf{tw}^{2})}$
This is bounded by $2^{\mathcal{O}(\textsf{tw}^{2})}\cdot|V(G)|^{\mathcal{O}(1)}$. This concludes the proof of the algorithmic result mentioned in Theorem 1. ∎
3.2 Conditional Lower Bound
In this section, we demonstrate that unless (ETH) fails, Labeled Contraction does not admit an algorithm running in time $2^{o(\textsf{tw}^{2})}\cdot|V(G)|^{\mathcal{O}(1)}$, where tw is the treewidth of the input graph $G$. To obtain the lower bound, we reduce from a special case of Vertex Cover on sub-cubic graphs, defined in [3] as follows:
Sub-Cubic Partitioned Vertex Cover (Sub-Cubic PVC) Input: A sub-cubic graph $G$; an integer $t$; for $i\in[t]$, an integer $k_{i}\geq 0$; a partition $\mathcal{P}=\{C_{1},\ldots,C_{t}\}$ of $V(G)$ such that $t\in\mathcal{O}(\sqrt{|V(G)|})$ and for all $i\in[t]$, $C_{i}$ is an independent set and $|C_{i}|\in\mathcal{O}(\sqrt{|V(G)|})$. Furthermore, for $\in[t], iji\neq j$, $|E(G[C_{i}\cup C_{j}])\cap E(G)|=1$. Question: Does $G$ have a vertex cover $X$ such that for all $i\in[t]$, $|X\cap C_{i}|\leq k_{i}$?
Proposition 3.1 (Theorem 3.9 in [3]).
Sub-Cubic PVC does not admit an algorithm running in time $2^{o(n)}$, unless the ETH fails.
We now prove the lower bound in Theorem 1.
Let $(G,\mathcal{P}=\{C_{1},C_{2},\ldots,C_{t}\},k_{1},\ldots,k_{t})$ be an instance of the Sub-Cubic PVC problem, where $n=|V(G)|$. We construct an instance $(G^{\prime},H^{\prime})$ of the Labeled Contraction problem as follows.
Construction of $G^{\prime}$: We begin by setting $V(G^{\prime})\supseteq V(G)$, i.e., all vertices of $G$ are copied into $G^{\prime}$ (without copying any edges). For each partition $C_{i}\in\mathcal{P}$, we introduce a set $S_{i}$ consisting of $\binom{|C_{i}|}{k_{i}+1}$ new vertices in $G^{\prime}$, where each vertex $s\in S_{i}$ corresponds uniquely to a subset $C_{i}^{\prime}\subseteq C_{i}$ of size exactly $k_{i}+1$. For each such vertex $s\in S_{i}$, we add edges between $s$ and every vertex in the subset $C_{i}$. Additionally, for every $C_{i}\in\mathcal{P}$, we introduce three auxiliary vertices $\{x_{i},y_{i},z_{i}\}$ in $G^{\prime}$. Each of these vertices is connected to every vertex in $C_{i}$. Moreover, we add edges $(z_{i},x_{i})$ and $(z_{i},y_{i})$. Each vertex $s\in S_{i}$ is connected to the corresponding vertex $x_{i}$. (Note that in the figure, an edge between $x_{i}$ and $S_{j}$ implies an edge between $x_{i}$ and every vertex in $S_{j}$.)
We now encode the edges in $G$. Consider a pair $i,j\in[t]$ and an edge $(u,v)\in E(G)$ with $u\in C_{i}$ and $v\in C_{j}$. By the problem definition, such an edge exists. We add edges $(u,x_{j})$ and $(v,x_{i})$ to $G^{\prime}$.
Finally, the set $\{y_{1},y_{2},\ldots,y_{t}\}$ is made into a clique by adding all possible edges between them, and for every $i,j\in[t]$, we add the edge $(x_{i},y_{j})$ in $G^{\prime}$.
Construction of $H^{\prime}$: We start constructing $H^{\prime}$ from $G^{\prime}$ by deleting all vertices in $G$ in it. Formally, $V(H^{\prime})=\left(\bigcup_{i=1}^{t}S_{i}\right)\cup\left(\bigcup_{i=1}^{t}\{x_{i},y_{i},z_{i}\}\right)$, that is, it includes all newly introduced vertices from the sets $S_{i}$ and the auxiliary vertices $x_{i},y_{i},z_{i}$ for each $i\in[t]$, but does not include any vertex from the original graph $G$. The edge set of $H^{\prime}$ is defined as follows. First, a clique is formed on the vertex set $\{y_{1},y_{2},\ldots,y_{t}\}\cup\{x_{1},x_{2},\ldots,x_{t}\}$, adding all edges between these vertices. Additionally, for every $i\in[t]$, we include the edges $(x_{i},z_{i})$ and $(y_{i},z_{i})$. Finally, for every $i\in[t]$ and each vertex $s\in S_{i}$, we add the edge $(s,y_{i})$.
This concludes the construction. We refer the reader to Figure 2 for an illustration of the reduction. We now prove the correctness of the reduction.
[…]
(a) Graph $G^{\prime}$
[…]
(b) Graph $H^{\prime}$
Figure 2: Illustration of the graphs $G^{\prime}$ and $H^{\prime}$ in the reduction for the lower bound of Theorem 1
$(\Rightarrow)$ Suppose we are given a solution to the Sub-Cubic PVC instance, that is, a vertex cover $X\subseteq V(G)$ such that for every $i\in[t]$, it holds that $|X\cap C_{i}|\leq k_{i}$. We now describe a contraction sequence on $G^{\prime}$ that yields $H^{\prime}$. For each partition $C_{i}\in\mathcal{P}$, we proceed as follows: For every vertex $a\in X\cap C_{i}$, we contract the edge $(a,x_{i})$. For every vertex $b\in C_{i}\setminus X$, we contract the edge $(b,y_{i})$.
Observe that since $X$ is a vertex cover of $G$, for each pair $i\neq j\in[t]$, there exists exactly one edge in $G$ between $C_{i}$ and $C_{j}$, say $(u,v)\in E(G)$, with $u\in C_{i}$ and $v\in C_{j}$. The contractions described above will result in $u$ being contracted to $x_{i}$ or $y_{i}$, and $v$ to $x_{j}$ or $y_{j}$. Since $u\in X$ or $v\in X$, at least one of the endpoints is contracted to an $x$-vertex, ensuring that each pair $(x_{i},x_{j})$ becomes adjacent in the resulting graph, thereby forming a clique on the set $\{x_{1},x_{2},\ldots,x_{t}\}$.
Furthermore, for each $i\in[t]$, since $|X\cap C_{i}|\leq k_{i}$, it follows that no subset of $C_{i}$ of size $k_{i}+1$ is entirely mapped to $x_{i}$. Therefore, for every vertex $s\in S_{i}$, which corresponds to a subset of $C_{i}$ of size $k_{i}+1$, at least one vertex in that subset is contracted to $y_{i}$. Consequently, the edge $(s,y_{i})$ appears in the resulting graph, as required by $H^{\prime}$. The edges $(x_{i},y_{i})$ and $(y_{i},z_{i})$ are preserved directly from the original construction. Thus, the contraction process yields graph $H^{\prime}$, completing the forward direction of the reduction.
$(\Leftarrow)$ Suppose we are given a solution to the Labeled Contraction instance, i.e., a sequence of edge contractions that transforms $G^{\prime}$ into $H^{\prime}$. We construct a corresponding solution $X\subseteq V(G)$ to the original Sub-Cubic PVC instance as follows.
Recall that the vertex set of $H^{\prime}$ is precisely $V(G^{\prime})\setminus V(G)$, which implies that every contracted edge in the solution involves a vertex from $V(G)$, more precisely vertices in $C_{i}$ for each $i\in[t]$. Furthermore, observe that for each $i\in[t]$, the vertices $z_{i}$ and those in $S_{i}$ do not share edges with each other or with vertices in $V(G)$ outside of $C_{i}$ neither in $G^{\prime}$ nor in $H^{\prime}$. In particular, the absence of edges of the form $(z_{i},s)$ for any $s\in S_{i}$, and the absence of edges $(z_{i},x_{j})$ or $(z_{i},y_{j})$ for $i\neq j$, implies that vertices in $C_{i}$ can only be either contracted to $x_{i}$ or $y_{i}$.
Now, for each vertex $s\in S_{i}$, we note that $s$ is adjacent to $y_{i}$ in $H^{\prime}$. In the construction of $G^{\prime}$, the vertex $s$ was made adjacent to a subset $C_{i}^{\prime}\subseteq C_{i}$ of size $k_{i}+1$. In order for $s$ to remain adjacent to $y_{i}$ in $H^{\prime}$, it is necessary that at least one vertex in $C_{i}^{\prime}$ is contracted to $y_{i}$. Thus, no subset of $C_{i}$ of size $k_{i}+1$ can be entirely contracted to $x_{i}$. This ensures that at most $k_{i}$ vertices from $C_{i}$ are contracted to $x_{i}$. Based on this observation, we define the vertex set $X\subseteq V(G)$ by including, for each $i\in[t]$, all vertices in $C_{i}$ that are contracted to $x_{i}$. From the argument above, we have $|X\cap C_{i}|\leq k_{i}$ for all $i\in[t]$.
It remains to show that $X$ is a vertex cover of $G$. Observe that in the construction of $G^{\prime}$, for each pair $\in[t], iji\neq j$, such that $E(G[C_{i}\cup C_{j}])\cap E(G)=\{(u,v)\}$ with $u\in C_{i}$, $v\in C_{j}$, we added edges $(u,x_{j})$ and $(v,x_{i})$ to $G^{\prime}$. Since $x_{i}$ and $x_{j}$ are adjacent in $H^{\prime}$, and since no edge directly connects $x_{i}$ and $x_{j}$ in $G^{\prime}$, at least one of the vertices $u$ or $v$ must be contracted to $x_{i}$ or $x_{j}$, respectively. This implies that either $u\in X$ or $v\in X$, and hence $X$ intersects every edge in $E(G)$. Therefore, $X$ is a valid vertex cover of $G$. This completes the proof of correctness for the reduction.
To upper bound the treewidth of the graph $G^{\prime}$, define the set $W=\{x_{i},y_{i},z_{i}\mid i\in[t]\}$. Observe that the graph $G^{\prime}-W$ consists of $t$ vertex-disjoint subgraphs, each corresponding to a complete bipartite graph between $C_{i}$ and $S_{i}$. For each $i\in[t]$, the set $C_{i}$ is of size $\mathcal{O}(\sqrt{n})$. This implies that the treewidth of each $G^{\prime}[C_{i}\cup S_{i}]$ is $\mathcal{O}(\sqrt{n})$. Therefore, treewidth of $G^{\prime}$ is at most $\textsf{tw}(G^{\prime})\leq|W|+\max_{i\in[t]}\textsf{tw}(G^{\prime}[C_{i}\cup S_{i}])=\mathcal{O}(\sqrt{n})$.
Now suppose, for the sake of contradiction, that there exists an algorithm $\mathcal{A}$ that solves Labeled Contraction in time $2^{o(\textsf{tw}(G^{\prime})^{2})}\cdot|V(G^{\prime})|^{\mathcal{O}(1)}$. We use $\mathcal{A}$ as a subroutine to construct an algorithm $\mathcal{B}$ for solving Sub-Cubic PVC. On input an instance $(G,\mathcal{P}=\{C_{1},C_{2},\ldots,C_{t}\},k_{1},\ldots,k_{t})$ of Sub-Cubic PVC with $n=|V(G)|$, algorithm $\mathcal{B}$ performs the following steps: It constructs the corresponding instance $(G^{\prime},H^{\prime})$ of Labeled Contraction as described in the reduction. This step takes time $2^{\mathcal{O}(\sqrt{n})}$. Then, it invoke algorithm $\mathcal{A}$ on the instance $(G^{\prime},H^{\prime})$ and return the result produced by $\mathcal{A}$.
The correctness of algorithm $\mathcal{B}$ follows directly from the correctness of the reduction. Since $\textsf{tw}(G^{\prime})=\mathcal{O}(\sqrt{n})$, the total running time of $\mathcal{B}$ is $2^{o(n)}$, which contradicts the ETH, by Proposition 3.1. This establishes the claimed conditional lower bound stated in Theorem 1.
4 $\mathsf{NP}$-hardness When Maximum Degree is Bounded
In this section, we prove Theorem 2 by establishing the $\mathsf{NP}$-hardness of the Labeled Contractibility problem. We restate it for reader’s conveience.
See 2
We present a polynomial-time reduction from a variation of the Positive-Not-All-3-SAT problem to Labeled Contractibility, even when both input graphs $G$ and $H$ have constant maximum degree. The source problem is formally defined as follows and is known to be $\mathsf{NP}$-complete [22].
Positive-Not-All-$(3,4)$-SAT Input: A $3$-CNF Boolean formula $\psi$ with a set of variables $\{x_{1},\dots,x_{n}\}$ and a set of clauses $\{C_{1},\dots,C_{m}\}$, where every literal is positive and each variable appears in at most 4 clauses. Question: Does there exist a truth assignment such that every clause is satisfied and, in each clause, not all three literals are assigned True?
For our reduction, we assume without loss of generality that the number of clauses $m$ is such that $2m+2$ is a power of 2. Let $p=2m+2=2^{q}$ for some integer $q$. We now describe the construction of two vertex-labeled graphs $G$ and $H$ from the formula $\psi$.
Construction of $H$: The target graph $H$ is a simple path on $p$ vertices, labeled $\{v_{1},v_{2},\ldots,v_{p}\}$.
Construction of $G$: We build $G$ by connecting two symmetric graph structures.
  1. 1.
    Path Structure: Create a path of $p$ vertices, labeled $v_{1},v_{2},\ldots,v_{p}$.
  2. 2.
    First Binary Tree: We build a complete binary tree of depth $q$ with its root at $v_{1}$. This tree has $2^{q}=p$ leaves.
    • The vertices on the shortest path from $v_{1}$ to the leaf corresponding to $v_{p/2}$ are labeled $v_{2},\ldots,v_{p/2-1}$.
    • The other leaves of the tree are labeled $\{C_{1},a_{1},C_{2},a_{2},\dots,C_{m},a_{m}\}$, alternating between clause vertices $C_{j}$ and auxiliary vertices $a_{j}$.
    • We add edges between consecutive auxiliary vertices, i.e., $(a_{i},a_{i+1})$ for all $i\in[m-1]$, creating a path among them.
  3. 3.
    Second Binary Tree: Construct a mirrored copy of the first binary tree, rooted at $v_{p}$. The leaves of this tree are labeled $\{C_{1}^{\prime},a_{1}^{\prime},C_{2}^{\prime},a_{2}^{\prime},\dots,C_{m}^{\prime},a_{m}^{\prime}\}$, corresponding to the unprimed vertices. The vertices on the shortest path from $v_{p}$ to $v_{p/2+1}$ are labeled $v_{p/2+2},\ldots,v_{p}$.
  4. 4.
    Connecting Structures: The two binary trees are connected via a central edge $(v_{p/2},v_{p/2+1})$. We also add an edge $(C_{j},C_{j}^{\prime})$ for each clause $j\in[m]$, as well as edges $(a_{1},v_{p/2})$ and $(a_{1}^{\prime},v_{p/2+1})$.
  5. 5.
    Variable Gadgets: For each variable $x_{i}\in\{x_{1},\dots,x_{n}\}$, we introduce a new vertex $u_{i}$. For every clause $C_{j}$ in which $x_{i}$ appears, we add edges connecting $u_{i}$ to the vertices $\{C_{j},a_{j},C_{j}^{\prime},a_{j}^{\prime}\}$.
This concludes the description of the reduction. Note that in the final graph $G$, each vertex has a maximum degree of at most 16. The internal structure of the tree and the connections are illustrated in Figure 3.
Next, we prove the correctness of the above reduction.
$(\Rightarrow)$ Let $\mathcal{W}=\{W_{1},W_{2},\dots,W_{p/2},W_{p/2+1},\cdots,W_{p}\}$ denote the witness structure corresponding to $H$. Let $\mathcal{W}(v_{i})=W_{i}$ for all $i\in[p]$, where $\mathcal{W}(u)$ is the witness set where $u\in V(H)$ is the representative vertex and $v_{i}\in W_{i}$. We will describe a sequence of contractions such that, for all $j\in[m]$, the vertices $\{C_{j},a_{j}\}$ are mapped to $W_{p/2}$ and the vertices $\{C_{j}^{\prime},a_{j}^{\prime}\}$ are mapped to $W_{p/2+1}$. More specifically, any vertex that is distance $i$ from $v_{1}$ must belong to the witness set $W_{i}$.
Suppose we are given a Yes-instance $\psi$ of Positive-Not-All-$(3,4)$-SAT, and construct $(G,H)$ as described above. Let $x_{i}$ be a variable assigned True in the satisfying assignment. For each clause $C_{j}$ in which $x_{i}$ appears, we contract the edge $(x_{i},a_{j})$. This contraction introduces the edge $(C_{j},a_{j})$, since both $C_{j}$ and $a_{j}$ are connected to the same vertex $x_{i}$. Similarly, for every clause $C_{j}$ in which a variable $x_{k}$ is assigned False, we contract the edge $(x_{k},a_{j}^{\prime})$, introducing the edge $(C_{j}^{\prime},a_{j}^{\prime})$.
Note that by the definition of Positive-Not-All-$(3,4)$-SAT, every clause must contain at least one variable set to True and at least one set to False. Therefore, for every clause $C_{j}$, there exists some $x_{i}$ with $x_{i}=\texttt{True}$ and $x_{i}\in C_{j}$, ensuring that $(C_{j},a_{j})$ appears in the contracted graph. Likewise, every $C_{j}^{\prime}$ will be associated with some $x_{k}=\texttt{False}$, ensuring that $(C_{j}^{\prime},a_{j}^{\prime})$ appears.
Further observe that now we have a path from $v_{i}\ \forall i\in[p]$ to all vertices that are at a $i$ distance from $v_{1}$. Contracting edges along this path will result in our desired witness structure.
($\Leftarrow$) Assume there exists a valid witness structure for a labeled contraction from $G$ to $H$. Note that witness sets are connected components and a set $W_{i}$ is adjacent only to $W_{i-1}$ and $W_{i+1}$. We use the following 2 observations: If a vertex $u$ is at a distance $\leq k$ from $v_{1}$, then it must belong to some $W_{i}$ for $i\in[1,k]$. Similarly any vertex at a distance $\leq(p-k)$ must belong to some $W_{j}$ such that $j\in[k,p]$. Therefore, we conclude that any vertex at the $k^{\text{th}}$ level (for $k\in[1,p/2]$) of the binary tree rooted at $v_{1}$ (respectively, at the $k^{\text{th}}$ level for $k\in[p/2+1,p]$ of the tree rooted at $v_{p}$) must belong to the witness set $W_{k}$.
Therefore, we see that to preserve the path structure of $H$, the only feasible witness sets for $\{C_{j},a_{j}\}$ and $\{C_{j}^{\prime},a_{j}^{\prime}\}$ are $W_{p/2}$ and $W_{p/2+1}$ respectively. For the other vertices note that we have a path in $G$ starting from $v_{i}\ \forall i\in\{1,2,\cdots,v_{p/2-1},v_{p/2+1},\cdots,v_{p}\}$ to all vertices that are at exactly $i$ distance from $v_{1}$. A similar construction can be only provided for the clauses by the variable vertices as any other contractions introduce edges across non-adjacent witness sets.
More specifically, recall that $H$ consists of a path on the vertex sequence $\{v_{1},v_{2},\dots,v_{p}\}$. In this path, the shortest path between $v_{p/2-1}$ and $v_{p/2+2}$ is exactly of length 3. In order to preserve this distance, any contraction sequence must not create shortcuts across this path, for instance, we cannot contract a vertex from the left subtree (rooted at $v_{1}$) into any vertex in the right subtree (rooted at $v_{p}$), or vice versa.
Now, for the edge $(C_{j},a_{j})$ to exist in $G$ before contraction, there must exist some $x_{i}$ such that the edge $(x_{i},a_{j})$ was contracted. This implies that $x_{i}$ was set to True in the assignment encoded by the contraction (since this is how $(C_{j},a_{j})$ is formed in the forward direction). Likewise, for the edge $(C_{j}^{\prime},a_{j}^{\prime})$ to appear, there must exist some $x_{k}$ set to False that was contracted to $a_{j}^{\prime}$.
Since the contraction sequence produces both $(C_{j},a_{j})$ and $(C_{j}^{\prime},a_{j}^{\prime})$, it must be the case that each clause $C_{j}$ has at least one true and at least one false literal—otherwise one of these edges could not be formed while preserving witness structure and distances.
Thus, the contraction sequence encodes a satisfying assignment of $\psi$ such that in every clause, not all literals are true. This shows that $\psi$ is a Yes-instance of Positive-Not-All-$3$-SAT.
This completes the proof for Theorem 2, that is Labeled Contraction is $\mathsf{NP}$-hard even when the maximum degree of both $G$ and $H$ are constant.
$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$v_{7}$$v_{8}$$C_{1}$$C_{1}^{\prime}$$a_{1}$$a_{1}^{\prime}$$C_{2}$$C_{2}^{\prime}$$a_{2}$$a_{2}^{\prime}$$C_{3}$$C_{3}^{\prime}$$a_{3}$$a_{3}^{\prime}$$x_{i}$
Figure 3: Example of the construction of $G$ with $m=3$. We only show one $x_{i}$ for convenience that appears in $C_{3}$
5 Parameterized by Solution Size Plus Degeneracy
We prove Theorem 3 and Theorem 4 in the following two subsections.
See 3 See 4
5.1 Conditional Lower Bound
We remark that most of the previous reduction about Labeled Contraction are from the variation of 3-SAT problem called Not-All-Equal-3-SAT. In this variation, the input is a $3$-CNF formula and the objective is to find an assignment that sets at least one but not all literals in every clause to set to True. We consider the following, different version of $3$-SAT as starting point of our reduction and present a simple proof from 3-SAT to get desired conditional lower bound.
11-in-33-SAT Input: A Boolean formula $\psi$ in 3-CNF Question: Does there exist a satisfying assignment such that exactly one literal is set to true in each clause?
Consider the following polynomial-time reduction from 3-SAT to 1-in-3-SAT where each clause $C^{i}=(x\lor\neg y\lor z)$ in the original formula is replaced by three clauses:
$C^{i}_{1}=(\neg x\lor a_{i}\lor b_{i}),\quad C^{i}_{2}=(y\lor b_{i}\lor c_{i}),\quad C^{i}_{3}=(\neg z\lor c_{i}\lor d_{i}),$
where $a_{i},b_{i},c_{i},d_{i}$ are fresh variables unique to the clause $C^{i}$. This reduction introduces a linear blow-up in the number of variables and clauses. Combining this reduction with the (ETH) [37] and the Sparsification Lemma [38], we obtain the following:
Proposition 5.1.
Unless the ETH fails, 11-in-33-SAT cannot be solved in time $2^{o(n+m)}$, where $n$ and $m$ are the number of variables and clauses in the input formula, respectively.
We now reduce 11-in-33-SAT to Labeled Contraction, ensuring that the number of vertices in $G$ and $H$ is linearly bounded by the number of variables and clauses in and instance of 11-in-33-SAT.
We now proceed with the proof of Theorem 3.
Proof.
Let $\psi$ be an instance of 11-in-33-SAT with variable set $\{x_{1},x_{2},\dots,x_{n}\}$ and clause set $\{C_{1},C_{2},\dots,C_{m}\}$. We construct a corresponding instance $(G,H)$ of Labeled Contraction such that $\psi$ is satisfiable if and only if $H$ is a labeled contraction of $G$.
We simultaneously define the graphs $G$ and $H$ as follows (See Figure 4 for an illustration.):
  • Global vertices: Add two special vertices $g_{T}$ and $g_{F}$ to both $G$ and $H$, connected by the edge $(g_{T},g_{F})$. These represent the ‘true’ and ‘false’ assignments, respectively, for literals.
  • Variable gadget: For each variable $x_{i}$, $i\in[n]$:
    • Add a vertex $u_{i}$ to both $G$ and $H$, representing variable $x_{i}$.
    • Add two vertices $v_{i}$ and $v^{\prime}_{i}$ to $G$, corresponding to literals $x_{i}$ and $\neg x_{i}$, respectively.
    • In $G$, connect $u_{i}$ to both $v_{i}$ and $v^{\prime}_{i}$. Also connect both $v_{i}$ and $v^{\prime}_{i}$ to $g_{T}$ and $g_{F}$.
    • In $H$, add edges $(u_{i},g_{T})$, $(u_{i},g_{F})$, and retain $(g_{T},g_{F})$.
  • Clause gadget: For each clause $C_{j}=(\ell_{j1}\lor\ell_{j2}\lor\ell_{j3})$, let $\mathsf{litv}(\ell)$ denote $v_{i}$ if $\ell=x_{i}$, and $v_{i}^{\prime}$ if $\ell=\neg x_{i}$.
    • Add four vertices $w_{j0},w_{j1},w_{j2},w_{j3}$ to both $G$ and $H$.
    • In $H$, connect $w_{j0}$ to $g_{T}$ and $g_{F}$.
    • In $G$, connect $w_{j0}$ to $\mathsf{litv}(\ell_{j1})$, $\mathsf{litv}(\ell_{j2})$, and $\mathsf{litv}(\ell_{j3})$.
    • For each $k\in\{1,2,3\}$, add edge $(w_{jk},g_{T})$ to both $G$ and $H$, and edge $(w_{jk},g_{F})$ only to $H$.
    • Finally, form a 6-cycle in $G$ by connecting the three $w_{jk}$’s with the three $\mathsf{litv}(\ell_{jk})$’s in an alternating cycle such that no two $w_{jk}$’s nor any two $\mathsf{litv}(\ell_{jk})$’s are adjacent.
This completes the reduction. It is easy to see that it takes polynomial time, and both graphs $G$ and $H$ have degeneracy at most $3$.
$g_{T}$$v_{i}$$u_{i}$$g_{F}$$v_{i}^{\prime}$
(a) Variable gadget in $G$
$g_{T}$$u_{i}$$g_{F}$
(b) Variable gadget in $H$
$g_{T}$$v_{i}$$v_{j}$$v_{k}^{\prime}$$w_{0}$$w_{1}$$w_{2}$$w_{3}$$g_{F}$
(c) Clause gadget in $G$
$g_{T}$$w_{0}$$w_{1}$$w_{2}$$w_{3}$$g_{F}$
(d) Clause gadget in $H$
Figure 4: Illustration of the reduction. (Top) Variable gadgets in $G$ and $H$ for a variable $(v_{i})$. (Bottom) Clause gadgets in $G$ and $H$ for some clause of the form $(v_{i}\lor v_{j}\lor v_{k}^{\prime})$.
We now show the equivalence between satisfiability of $\psi$ and labeled contractibility of $G$ into $H$.
($\Rightarrow$) Suppose there exists a satisfying assignment $\alpha:\{x_{1},\dots,x_{n}\}\to\{\texttt{True},\texttt{False}\}$ such that exactly one literal is true in each clause. Define the contraction sequence $S$ as follows:
  • For each $x_{i}$, if $\alpha(x_{i})=\texttt{True}$, then add edges $(v_{i},g_{T})$, $(v_{i}^{\prime},g_{F})$ to $S$; and if $\alpha(x_{i})=\texttt{False}$, then add edges $(v_{i},g_{F})$, $(v_{i}^{\prime},g_{T})$ to $S$.
This ensures the correct neighborhood for $u_{i}$ in $H$, as it will have edges to both $g_{T}$ and $g_{F}$.
Now consider any clause $C_{j}$. Since exactly one literal is true under $\alpha$, exactly one $\mathsf{litv}(\ell_{jk})$ is contracted into $g_{T}$, and the remaining two are contracted into $g_{F}$. The contraction of the $w_{jk}$-triangle via the 6-cycle then ensures:
  • The edge $(w_{j0},g_{T})$ is introduced by the literal assigned true;
  • The edge $(w_{j0},g_{F})$ is introduced by the literals assigned false;
  • All three edges $(g_{F},w_{j1}),(g_{F},w_{j2}),(g_{F},w_{j3})$ are created.
Hence, the resulting graph after the contraction is precisely $H$.
($\Leftarrow$) Conversely, suppose $H$ is a labeled contraction of $G$. We argue that this implies a satisfying assignment for $\psi$.
The only way to create both edges $(u_{i},g_{T})$ and $(u_{i},g_{F})$ in $H$ is by contracting $v_{i}$ and $v_{i}^{\prime}$ into $g_{T}$ and $g_{F}$, respectively, or vice versa. This naturally defines a valid assignment:
[…]
Next, observe that to produce all three edges $(g_{F},w_{j1}),(g_{F},w_{j2}),(g_{F},w_{j3})$, at least two of the literals in clause $C_{j}$ must contract to $g_{F}$, while one contracts to $g_{T}$. Thus, in each clause, exactly one literal must be set to true under $\alpha$. Therefore, $\alpha$ satisfies $\psi$ as a 1-in-3-SAT assignment.
Each vertex in the variable and clause gadgets has degree at most 3 in $G$ and at most 2 in $H$, so the degeneracy of both graphs is bounded by a constant. The size of $G$ is linear in $n+m$, and the construction can be completed in polynomial time. Hence, if Labeled Contraction can be solved in time $2^{o(|V(G)|+|E(G)|)}$, then 1-in-3-SAT can be solved in time $2^{o(n+m)}$, which contradicts ETH according to Proposition 5.1. This completes the proof of Theorem 3. ∎
5.2 Algorithmic Result
We first present proof of Theorem 4 establishing an alternate $\mathsf{FPT}$ algorithm and later mention its consequences. Recall that the theorem states that the Labeled Contraction problem admits an algorithm with running time $(\delta(H)+1)^{k}\cdot|V(G)|^{\mathcal{O}(1)}$, where $k=|V(G)\setminus V(H)|$.
Let $(G,H)$ be an instance of the Labeled Contraction problem. With sanity check, we can assume that $V(H)\subseteq V(G)$.
The goal is to decide whether there exists a valid witness structure $\mathcal{W}=\{W_{h}\subseteq V(G)\mid h\in V(H)\}$ such that contracting each $W_{h}$ into the vertex $h$ yields the graph $H$.
Let $k=|V(G)\setminus V(H)|$ denote the number of vertices in $G$ that are not in $H$, and hence must be contracted. Each such vertex $x\in V(G)\setminus V(H)$ must be assigned to exactly one vertex $h\in V(H)$, corresponding to the witness set $W_{h}$ into which $x$ will be contracted.
Define a mapping $\phi:V(G)\to V(H)$ as follows: $(i)$ For each $v\in V(H)$, set $\phi(v)=v$; $(ii)$ For each $x\in V(G)\setminus V(H)$, $\phi(x)$ denotes the vertex in $V(H)$ to which $x$ is assigned (i.e., the representative vertex of the witness set containing $x$). A mapping $\phi$ defines a candidate witness structure, and we wish to determine whether it corresponds to a valid contraction of $G$ into $H$.
To bound the number of candidate functions $\phi$, we exploit structural properties of the graph $H$. Let $\delta(H)$ denote the degeneracy of $H$. By definition, $H$ admits an ordering $v_{1},v_{2},\ldots,v_{n}$ such that each vertex has at most $\delta(H)$ neighbors among its predecessors. This implies that $H$ is $(\delta(H)+1)$-colorable, and such a coloring can be computed in polynomial time.
Fix a proper coloring $c:V(H)\to[\delta(H)+1]$. For each vertex $x\in V(G)\setminus V(H)$, let $N_{H}(x):=\{v\in V(H)\mid\{x,v\}\in E(G)\}$. If two vertices $y_{1},y_{2}\in N_{H}(x)$ belong to the same color class (i.e., $c(y_{1})=c(y_{2})$) and $\{y_{1},y_{2}\}\notin E(H)$, then assigning $x$ to either $y_{1}$ or $y_{2}$ would result in the edge $\{y_{1},y_{2}\}$ being introduced during the contraction process—contradicting the assumption that $H$ is the target graph. Therefore, for each $x\in V(G)\setminus V(H)$, there are at most $\delta(H)+1$ valid choices for $\phi(x)$. These choices corresponds at most one vertex in each color class. Thus, the total number of candidate assignments $\phi$ is bounded by $(\delta(H)+1)^{k}$. For each such assignment, we can verify in polynomial time whether the induced contraction yields $H$, by simulating the contraction and checking edge and label consistency. This yields an algorithm with running time $(\delta(H)+1)^{k}\cdot|V(G)|^{\mathcal{O}(1)}$. This completes the proof for Theorem 4
The heuristic proper coloring in the proof can be replaced by by optimal proper coloring to obtain the following corollary.
Corollary 1.
The Labeled Contraction problem admits an algorithm with running time $2^{|V(H)|}\cdot|V(H)|^{\mathcal{O}(1)}+\chi(H)^{k}\cdot|V(G)|^{\mathcal{O}(1)}$, where $\chi(H)$ is the chromatic number of $H$.
Proof.
Let $\chi(H)$ denote the chromatic number of $H$. Clearly, $\chi(H)\leq\delta(H)+1$. Fix an optimal coloring of $H$ with $\chi(H)$ colors. As before, the coloring constrains the valid targets for each vertex $x\in V(G)\setminus V(H)$ to at most $\chi(H)$ choices, since assigning $x$ to a vertex $h\in V(H)$ may introduce forbidden adjacency depending on its neighbors. Thus, the number of candidate assignments is bounded by $\chi(H)^{k}$, and for each such assignment, validity can be checked in polynomial time. Recall that an optimal coloring of $H$ can be found in time $2^{\mathcal{O}(|V(H)|)}$. This yields an algorithm with overall running time as mentioned. ∎
[41, Lemma 6] states that if $H$ can be obtained from a graph $G$ by at most $k$ edge contractions, then $\delta(H)\leq\delta(G)+k$. Motivated by this result, we present an alternative degree bound that depends on the size of the graph.
Lemma 5.1.
If $H$ can be obtained from $G$ by $k$ edge contractions, then
$\delta(H)\leq\delta(G)\cdot\frac{2|V(G)|}{|V(G)|-k}.$
Proof.
We use the following standard inequalities. For graph $G$ with degeneracy $\delta(G)$, it holds that $|E(G)|\leq\delta(G)\cdot|V(G)|$. This is an upper bound on the number of edges. Further for any graph $H$, $\frac{1}{2}\cdot\delta(H)\cdot|V(H)|\leq|E(H)|$. This follows from the fact that $\delta(H)$ is a lower bound on the average degree. Since, edge contraction does not increase the number of edges, we have $|E(H)|\leq|E(G)|$. Furthermore, the number of vertices in $H$ is |V(H)|=|V(G)|k|V(H)|=|V(G)|-k. Combining the bounds, we have: $\frac{1}{2}\cdot\delta(H)\cdot(|V(G)|-k)\leq|E(H)|\leq\delta(G)\cdot|V(G)|.$ Solving for $\delta(H)$ gives $\delta(H)\leq(\delta(G)\cdot 2|V(G)|)/({|V(G)|-k})$. ∎
In particular, the above lemma implies that if $|V(G)|\geq(1+\varepsilon)\cdot k$ for some $\varepsilon>0$, then $\delta(H)\leq c_{\varepsilon}\cdot\delta(G)$ for some constant $c_{\varepsilon}$.
6 Brute-force Algorithm and its Optimality
We now prove Theorem 5. We restate it for reader’s conveience.
See 5
Lemma 6.1.
The Maximum Common Labeled Contraction problem admits an algorithm with running time $|V(H)|^{\mathcal{O}(|V(G)|)}$.
Proof.
We begin by describing an algorithm for the Labeled Contractibility problem with running time $|V(H)|^{\mathcal{O}(|V(G)|)}$, which we later use as a subroutine to solve Maximum Common Labeled Contraction.
As established in Observation 2 of [41], a graph $H$ is a labeled contraction of a graph $G$ if and only if there exists a witness structure $W=\{W_{1},\ldots,W_{|V(H)|}\}$ of $G$ into $H$. Recall that a witness structure is a partition of $V(G)$ into $|V(H)|$ non-empty sets, where each $W_{i}$ corresponds to a distinct vertex of $H$.
The algorithm exhaustively enumerates all such candidate partitions and checks their validity. Formally, it distributes all the vertices in $V(G)\setminus V(H)$ into the $|V(H)|$ parts. Since each of the |V(G)||V(H)||V(G)|-|V(H)| vertices can be assigned to any of the $|V(H)|$ parts, the number of such candidate partitions is at most $(|V(H)|)^{|V(G)|-|V(H)|}$.
For each candidate witness structure $W$, the algorithm performs the following checks: $(i)$ For each $W_{j}\in W$, verify that the induced subgraph $G[W_{j}]$ is connected. $(ii)$ For all distinct $u,v\in V(H)$, verify that $uv\in E(H)$ if and only if there exists at least one edge in $G$ between some vertex in $W(u)$ and some vertex in $W(v)$. Both these steps take time which is polynomial in $|V(G)|$. If a partition satisfies both conditions, then $W$ is a valid witness structure certifying that $H$ is a labeled contraction of $G$. If no such structure is found, the algorithm correctly concludes that $H$ is not a labeled contraction of $G$. Thus, the total running time of this procedure is bounded by $(|V(H)|)^{|V(G)|-|V(H)|}\cdot\mathcal{O}(|V(G)|^{2})=|V(H)|^{\mathcal{O}(|V(G)|)}$.
We now describe the algorithm for Maximum Common Labeled Contraction. Let $G$ and $H$ be the two input graphs. If a graph $H^{\prime}$ is a common labeled contraction of both $G$ and $H$, then $H^{\prime}$ must be a labeled contraction of both. The algorithm enumerates all possible candidate graphs $H^{\prime}$ that can be obtained as a labeled contraction of $H$, and for each candidate, checks whether it is also a labeled contraction of $G$.
To bound the number of such candidate graphs $H^{\prime}$, we count the number of valid witness structures from $H$ to $H^{\prime}$. For a fixed size $n^{\prime}=|V(H^{\prime})|$, there are $\binom{|V(H)|}{n^{\prime}}$ ways to choose the representative vertices of $H^{\prime}$, and the remaining $|V(H)|-n^{\prime}$ vertices can be assigned to these parts in $(n^{\prime})^{|V(H)|-n^{\prime}}$ ways. Summing over all possible values of $n^{\prime}$, the total number of such graphs is bounded by $\sum_{n^{\prime}=1}^{|V(H)|}\binom{|V(H)|}{n^{\prime}}\cdot(n^{\prime})^{|V(H)|-n}\leq|V(H)|^{|V(H)|}\cdot 2^{|V(H)|}=|V(H)|^{\mathcal{O}(|V(H)|)}$.
For each such candidate $H^{\prime}$, we check: $(i)$ whether $H^{\prime}$ is a labeled contraction of $H$, which takes time $|V(H)|^{\mathcal{O}(|V(H)|)}$, and $(ii)$ whether $H^{\prime}$ is a labeled contraction of $G$, using the subroutine above in time $|V(H^{\prime})|^{\mathcal{O}(|V(G)|)}\leq|V(H)|^{\mathcal{O}(|V(G)|)}$. If both checks succeed and the total number of contractions used in both $G$ and $H$ is at most $k$, the algorithm concludes that it is a Yes instance. Hence, the overall running time of the algorithm is $|V(H)|^{\mathcal{O}(|V(G)|)}$. This concludes the proof of the lemma. ∎
Now we present the conditional lower bound, demonstrating that the algorithm is essentially optimal unless the Exponential Time Hypothesis (ETH) fails.
Our lower bound is established through a polynomial-time reduction from the Cross Matching problem, introduced by Fomin et al. [28]. This problem has been a hardness source for several other contraction-based graph problems.
Cross Matching Input: A graph $G^{\prime}$ together with a partition $(A,B)$ of $V(G^{\prime})$ such that $|A|=|B|$. Question: Does there exist a perfect matching $M\subseteq E(G^{\prime})$ such that each edge in $M$ connects a vertex from $A$ to one from $B$, and the graph $G^{\prime}/M$ is a clique?
Proposition 6.1 (Lemma 4.1 in [28]).
Unless the ETH fails, Cross Matching cannot be solved in time $n^{o(n)}$, where $n=|A|=|B|$.
We now present a polynomial-time reduction from Cross Matching to Labeled Contractibility.
[…]
(a) Graph $G$
[…]
(b) Graph $H$
Figure 5: Illustration of the graphs $G$ and $H$ in the reduction. The required adjacency across $\{x_{1},x_{2}\}$ and AA forces a candidate solution to form a matching across AA and $B$.
Lemma 6.2.
Unless the ETH fails, the Labeled Contractibility problem does not admit an algorithm running in time $|V(H)|^{o(|V(G)|)}$.
Proof.
Let $(G^{\prime},A,B)$ be an instance of Cross Matching with $|A|=|B|=n$. We construct an equivalent instance $(G,H)$ of Labeled Contractibility as follows.
The graph $G$ is obtained from $G^{\prime}$ by introducing two new vertices, $x_{1}$ and $x_{2}$, and adding edges from each to all vertices in $B$. Formally, $V(G)=V(G^{\prime})\cup\{x_{1},x_{2}\}$ and $E(G)=E(G^{\prime})\cup\{(x_{1},b),(x_{2},b)\mid b\in B\}$.
The target graph $H$ is defined on the vertex set $V(H)=A\cup\{x_{1},x_{2}\}$, with the following edge set: (i) all pairs $\{a_{1},a_{2}\}\subseteq A$ are adjacent (i.e., $A$ induces a clique); and (ii) both $x_{1}$ and $x_{2}$ are adjacent to every vertex in $A$. Vertices $x_{1}$ and $x_{2}$ are not adjacent. Equivalently, $H$ is a complete graph on $A\cup\{x_{1},x_{2}\}$ minus the edge $(x_{1},x_{2})$. See Figure 5 for an illustration.
We prove that $(G^{\prime},A,B)$ is a Yes-instance of Cross Matching if and only if $(G,H)$ is a Yes-instance of Labeled Contractibility.
($\Rightarrow$) Assume $(G^{\prime},A,B)$ is a Yes-instance. Then there exists a perfect matching $M$ between $A$ and $B$ such that the contracted graph $G^{\prime}/M$ is a clique. We construct a labeled contraction sequence $S$ on $G$ by contracting each edge $(a,b)\in M$, with $a\in A$ and $b\in B$, in any order. Since $M$ is a matching, all contracted edges are vertex-disjoint.
After all contractions in $S$, each vertex $b\in B$ is removed, and the vertex set of $G/S$ is $A\cup\{x_{1},x_{2}\}$. The subgraph induced by $A$ in $G/S$ is a clique, because $G^{\prime}/M$ is a clique and the contractions only merge vertices from $B$ into $A$. Furthermore, since $x_{1}$ and $x_{2}$ were connected to every $b\in B$ in the original graph, and each $b$ was contracted into some $a\in A$, these contractions introduce edges from both $x_{1}$ and $x_{2}$ to every $a\in A$. However, since $x_{1}$ and $x_{2}$ were not adjacent in the original graph and were not part of any contraction, the edge $(x_{1},x_{2})$ is not introduced. The resulting graph $G/S$ is precisely $H$.
($\Leftarrow$) Assume $(G,H)$ is a Yes-instance, meaning there exists a sequence $S$ of labeled contractions such that $G/S=H$. By construction, $V(G)=A\cup B\cup\{x_{1},x_{2}\}$ and $V(H)=A\cup\{x_{1},x_{2}\}$, which implies that all vertices in $B$ must be removed during the contraction sequence. In the labeled contraction model, a vertex is removed only when it is the second vertex in a contracted pair $(u,v)$. Thus, each $b\in B$ must appear as the second vertex in some contraction $(u,b)\in S$.
We now show that each $b\in B$ must be contracted into a distinct vertex $a\in A$. Suppose for contradiction that some $a\in A$ does not serve as the representative of any contracted pair. Consider the edge $(a,x_{1})$ in $H$. Since $a$ and $x_{1}$ are not adjacent in the original graph $G$, this edge must have been created by a contraction involving a common neighbor. The only common neighbors of $a$ and $x_{1}$ in $G$ are vertices in $B$. To create an edge $(a,x_{1})$, a vertex $b\in B$ must be contracted into either $a$ or $x_{1}$.
Case 1: $b$ is contracted into $a$. This creates the edge $(a,x_{1})$. Since $b$ is also a neighbor of $x_{2}$, this contraction also creates an edge $(a,x_{2})$, which is present in $H$. Case 2: $b$ is contracted into $x_{1}$. This creates an edge $(x_{1},a)$. However, since $b$ is also a neighbor of $x_{2}$, this contraction would also create an edge $(x_{1},x_{2})$, which is a contradiction since $(x_{1},x_{2})\notin E(H)$. Therefore, for each $b\in B$, there must be a contraction that merges it into some vertex in $A$. Since $|A|=|B|=n$, this implies that each vertex in $B$ is contracted into a distinct vertex in $A$, which forms a perfect matching $M$ between $A$ and $B$.
Finally, we argue that $G^{\prime}/M$ is a clique. Since $G/S=H$, and the subgraph induced by $A$ in $H$ is a clique, the same must hold in $G^{\prime}/M$. This is because the contractions only merge vertices from $B$ into $A$, preserving the structure among the vertices of $A$. Thus, $(G^{\prime},A,B)$ is a Yes-instance of Cross Matching.
The construction of $(G,H)$ from $(G^{\prime},A,B)$ takes polynomial time. The sizes of the graphs are $|V(G)|=2n+2$ and $|V(H)|=n+2$. Therefore, if Labeled Contractibility could be solved in time $|V(H)|^{o(|V(G)|)}$, then Cross Matching could be solved in time $n^{o(n)}$, which contradicts Proposition 6.1. This completes the proof of the lemma. ∎
7 Conclusion
In this article, we advanced the study of the parameterized complexity of Labeled Contractibility, building upon the foundational work of Lafond and Marchand [41]. We developed a constructive dynamic programming algorithm with a running time of $2^{\mathcal{O}(\textsf{tw}^{2})}\cdot n^{\mathcal{O}(1)}$, which is a significant improvement over the previously known non-constructive meta-algorithmic approach. Furthermore, we complemented this result with a tight, matching conditional lower bound under the Exponential Time Hypothesis (ETH), which is a relatively rare phenomenon. Our work resolves several open questions from [41]. We provided an improved $\mathsf{FPT}$ algorithm with respect to the combined parameter of solution size and degeneracy. We also established that the straightforward brute-force approach for the problem is optimal under ETH, thereby providing a complete complexity landscape for this particular perspective.
We conclude by highlighting some directions for future research. We conjecture that our dynamic programming framework over tree decompositions could be extended to solve the more general Maximum Common Labeled Contraction problem. Additionally, it would be interesting to explore whether fixed-parameter algorithms with better running time can be obtained when parameterized by the feedback vertex set number or the vertex cover number.

### v2

@@
 This equivalence allows complexity results to be transferred between the two problems.
-The same work provides additional applications and results.
+We refer readers to the same work [41] for additional results and applications.
 Lafond and Marchand [41] showed that Labeled Contractibility is ${\mathsf{W}}[1]$-hard when parameterized by the natural parameter $k=|V(G)\setminus V(H)|$, the number of allowed edge contractions.
@@
 The resulting graph is denoted by $G/(u,v)$.
+The same is illustrated on a toy example in Figure 1.
+GG12534MM1254
+Figure 1: Contraction along edge $(2,3)$
 Note that in general, $G/(u,v)\neq G/(v,u)$, as the vertex that is retained after the contraction (and thus the label) is different.
@@
   • For any distinct $u,v\in V(H)$, the edge $(u,v)\in E(H)$ if and only if the sets in $\mathcal{W}$ containing $u$ and $v$ are adjacent in $G$.
+We illustrate the same in Figure 2.
+GG1253467HH1237
+Figure 2: Shaded regions in $G$ form the witness sets, each containing exactly one representative vertex (circled). Contracting each witness set yields $H$.
 There is a natural bijection between the sets in $\mathcal{W}$ and the vertices of $H$ via their representatives.
@@
   • A forget node $t$ has one child $t_{1}$ such that $X_{t}=X_{t_{1}}\setminus\{x\}$, where $x\in X_{t_{1}}$.
-  • An introduce edge node $t$ has one child $t^{\prime}$ with $X_{t}=X_{t^{\prime}}$, and is labeled with an edge $\{u,v\}\in E(G)$ such that $\{u,v\}\subseteq X_{t}$.
-We say that the edge $\{u,v\}$ is introduced at $t$;
+  • An introduce edge node $t$ has one child $t^{\prime}$ with $X_{t}=X_{t^{\prime}}$, and is labeled with an edge $(u,v)\in E(G)$ such that $\{u,v\}\subseteq X_{t}$.
+We say edge $(u,v)$ is introduced at $t$.
   • A join node $t$ has two children $t_{1}$ and $t_{2}$ such that $X_{t}=X_{t_{1}}=X_{t_{2}}$.
@@
 This choice is crucial because it ensures that for any edge $(u,v)\in E(H)$, there exists some bag $X_{t}$ such that $u,v\in X_{t}$, a condition that is not guaranteed in a tree decomposition of $G$ alone.
-For each node $t\in\mathcal{T}$, we define $G_{t}$ and $H_{t}$ to be the subgraphs of $G$ and $H$, respectively, induced by the union of all bags in the subtree of $\mathcal{T}$ rooted at $t$.
-We now introduce the key definitions for the DP-state at a node $t$ of the tree-decomposition to convey the main ideas of the algorithm. The formal proof follows.
-3.1.1 Witness Structures and Functional Representation
-Assume that $H$ is a contraction of $G$, and let $\mathcal{W}=\{W_{1},\dots,W_{|V(H)|}\}$ denote a witness structure for this contraction.
-We represent $\mathcal{W}$ functionally using a pair of auxiliary functions: a parent function and an originator function.
-Definition 3.3 (Parent function).
-Given a witness structure $\mathcal{W}$, a function $\sigma^{\prime}:V(G)\to V(G)$ is a parent function with respect to $\mathcal{W}$ if there exists a collection of rooted trees $\{T_{i}\}_{i=1}^{k}$, where each $T_{i}$ is a spanning tree of $G[W_{i}]$ rooted at the unique vertex of $W_{i}\cap V(H)$, such that:
-  1. 1.
-    For all $u\in W_{i}\setminus V(H)$, $\sigma^{\prime}(u)$ is the parent of $u$ in $T_{i}$; and
-  2. 2.
-    For the unique vertex $u\in W_{i}\cap V(H)$, $\sigma^{\prime}(u)=u$.
-Definition 3.4 (Originator function).
-Given a witness structure $\mathcal{W}$, a function $\tau^{\prime}:V(G)\to V(H)$ is an originator function with respect to $\mathcal{W}$ if for every $u\in W_{i}$, $\tau^{\prime}(u)$ is the unique vertex in $W_{i}\cap V(H)$.
-Note that any valid originator function $\tau^{\prime}$ defines a partition of $V(G)$ into the witness sets $\mathcal{W}$.
-Also, for any vertex $u\in V(G)$, there exists a bag that contains both $u$ and $\sigma^{\prime}(u)$, but not necessarily $\tau^{\prime}(u)$.
-A pair $(\sigma^{\prime},\tau^{\prime})$ is said to be consistent if for every $u\in V(G)$ $\tau^{\prime}(u)=\tau^{\prime}(\sigma^{\prime}(u))$, i.e.
-$u$ and its parent $\sigma^{\prime}(u)$ belong to the same witness set.
-In the remainder of the discussion, we assume all parent-originator pairs to be consistent unless stated otherwise.
-Although $\tau^{\prime}$ can be derived from $\sigma^{\prime}$ by following parent pointers up to a vertex in $V(H)$, we maintain both functions explicitly for algorithmic clarity.
-Definition 3.5 (Witness impression).
-Every witness structure $\mathcal{W}$ corresponds to pair $(\sigma^{\prime},\tau^{\prime})$, called its witness-impression, where $\sigma^{\prime}$ is a parent function and $\tau^{\prime}$ is the corresponding originator function.
-Definition 3.6 (Seed impression).
-Let $(\sigma^{\prime},\tau^{\prime})$ be an impression of a witness structure.
-A pair of partial functions $(\sigma,\tau)$, with $\sigma:V(G_{t})\to V(G_{t})\cup\{\star\}$ and $\tau:V(G_{t})\to V(H_{t})\cup\{\star\}$, is a seed impression of $(\sigma^{\prime},\tau^{\prime})$ at $t$ if:
-  • For every vertex $u\in V(G_{t})\setminus X_{t}$, $\sigma(u)=\sigma^{\prime}(u)$ and $\tau(u)=\tau^{\prime}(u)$.
-  • For a vertex $u\in X_{t}$, if $\sigma^{\prime}(u)\in V(G_{t})$ then $\sigma(u)=\sigma^{\prime}(u)$ otherwise $\sigma(u)=\star$.
-The function $\tau$ is derived from $\sigma$ by following parent pointers up to either a vertex in $V(H)$ or $\star$.
-Note that $(\sigma,\tau)$ is not merely the restriction of $(\sigma^{\prime},\tau^{\prime})$ to $V(G_{t})$, since a vertex $u$ such that $\sigma^{\prime}(u)\notin V(G_{t})$ might satisfy $\tau^{\prime}(u)\in V(G_{t})$, but we still assign $\tau(u)=\star$.
-Note that in Figure 1, $\tau^{\prime}(u_{1})\in G_{t}$ but since $\sigma^{\prime}(u_{1})\notin G_{t}$ we assign $\tau(u_{1})=\star$.
-We now classify each vertex in the bag $X_{t}$ into one of three types, according to the seed impression $(\sigma,\tau)$:
-Definition 3.7 (Types of vertices).
-Let $t\in V(\mathcal{T})$ be a node with bag $X_{t}$, and let $(\sigma,\tau)$ be a seed impression at $t$. Then each vertex $u\in X_{t}$ is assigned one of the following types:
-  • Requester: $\sigma(u)\in V(G_{t})$ and $\tau(u)\in V(H_{t})$ (i.e., both are defined); $u$ is ready to request adjacency requirements in $H$.
-  • Dependent: $\sigma(u)\in V(G_{t})$ but $\tau(u)=\star$; $u$ has been contracted into its parent and depends on the parent node’s decision regarding the partition.
-  • Explorer: $\sigma(u)=\star$ (implying $\tau(u)=\star$); $u$ is still exploring who its parent will be.
-We classify the pair of requester vertices into the following three parts.
-Definition 3.8 (Requester pair classification).
-Let $u,v\in X_{t}$ be requester vertices with $\tau(u)=h_{u}$, $\tau(v)=h_{v}$. We say the pair $(u,v)$ is
-  • invalid (inv) if $(h_{u},h_{v})\notin E(H)$ but $\tau^{-1}(h_{u}),\tau^{-1}(h_{v})$ are adjacent in $G_{t}$,
-  • unsatisfied (unsat) if $(h_{u},h_{v})\in E(H)$ and $\tau^{-1}(h_{u}),\tau^{-1}(h_{v})$ are not adjacent in $G_{t}$, and
-  • satisfied (sat), otherwise.
-For a vertex $u\in V(G)$, we define the set $\textsf{des}(u)\subseteq V(G)$ of descendants of $u$ with respect to a function $\sigma^{\prime}:V(G)\to V(G)$ as follows:
-$\textsf{des}(u)=\left\{v\in V(G)\;\middle|\;\text{there exists an integer }\ell\geq 0\text{ such that }\sigma^{\prime(\ell)}(v)=u\right\},$
-where $\sigma^{\prime(0)}(v)=v$, and for $\ell\geq 1$, $\sigma^{\prime(\ell)}(v)=\sigma^{\prime}(\sigma^{\prime(\ell-1)}(v))$.
-That is, $\textsf{des}(u)$ consists of all vertices $v$ for which repeated applications of $\sigma^{\prime}$ eventually reach $u$, including $u$ itself.
-Definition 3.9 (Pseudo-adjacency).
-Let $u,v$ be two vertices in $X_{t}$. Then, we say $u$ and $v$ are pseudo-adjacent if sets $\textsf{des}(u)$ and $\textsf{des}(v)$ are adjacent in $G_{t}$.
-3.1.2 Signatures and Dynamic Programming States
-Suppose the graph $G$ admits an $H$-witness structure $\mathcal{W}$, and let $(\sigma^{\prime},\tau^{\prime})$ be the corresponding witness-impression as specified in Definition 3.5.
-Let $T$ be a tree decomposition of the graph $G\cup H$, and let $t$ be a node of $T$.
-Definition 3.10 (Partial Solution).
-We define a partial solution at node $t$ as a tuple $P=(\sigma,\tau,\mathcal{R},\mathcal{A}),$ with the following components:
-  • $(\sigma,\tau)$ is a seed impression of the witness-impression $(\sigma^{\prime},\tau^{\prime})$, restricted to the bag $X_{t}$ at $t$, as specified in Definition 3.6,
-  • $\mathcal{R}\subseteq R\times R\times\{\textsf{inv},\textsf{sat},\textsf{unsat}\}$ is a set of labeled pairs, where $R\subseteq X_{t}$ denotes the set of requester vertices contained in $X_{t}$ as per Definition 3.7.
-Each element $(u,v,\ell)\in\mathcal{R}$ represents a label $\ell\in\{\textsf{inv},\textsf{sat},\textsf{unsat}\}$ assigned to the pair $(u,v)$, as per Definition 3.8,
-  • $\mathcal{A}\subseteq X_{t}\times X_{t}$ records unordered pairs of vertices in $X_{t}$ that are pseudo-adjacent to each other, as per Definition 3.9.
-We now refer to Figure 1 to provide intuition for the above definitions.
-The figure shows the graph $H$ together with a part of the tree decomposition of $G\cup H$ (we do not explicitly draw $G$).
-The rounded outlines indicate the witness structure for the contraction, whenever it exists: each group contains exactly one node of $H$, representing the collection of vertices that contract to it.
-Note that $\sigma(u_{2})=u_{1}$, $\tau(u_{2})=\star$, and $\tau(u_{1})=\sigma(u_{1})=\star$.
-Since the edge $(u_{3},u_{4})$ is an edge we processed earlier in the DP, we may store $(u_{7},u_{5},\textsf{sat})\in\mathcal{R}$ for $X_{t}$.
-On the other hand, because the edge $(u_{6},u_{5})$ has not yet been introduced, the edge $(h_{2},h_{3})$ does not appear in the current structure, so we have $(u_{2},u_{5},\textsf{unsat})$.
-Furthermore, observe that $(u_{8},u_{5})$ and $(u_{7},u_{5})$ belong to $\mathcal{A}$, whereas $(u_{9},u_{5})\notin\mathcal{A}$.
-In this way, all relevant sets are defined for the dynamic programming.
-$u_{8}$$u_{7}$$u_{1}$$h_{2}$$u_{2}$$u_{9}$$u_{5}$$u_{3}$$h_{1}$$u_{4}$$h_{3}$$u_{6}$ $X_{t}$$G_{t}$RestofGRest~of~G
+For each node $t\in\mathcal{T}$, we define $G_{t}$ and $H_{t}$ to be the subgraphs of $G$ and $H$, respectively, induced by the union of all bags in the sub-tree of $\mathcal{T}$ rooted at $t$.
+$u_{8}$$u_{7}$$u_{1}$$h_{2}$$u_{2}$$u_{9}$$u_{5}$$u_{3}$$h_{1}$$u_{4}$$h_{3}$$u_{6}$ $W_{1}$ $W_{2}$ $W_{3}$$X_{t}$$G_{t}$RestofGRest~of~G
 (a) Illustration of $G_{t}$ and $X_{t}$
@@
 Figure 3: Depicting definitions relevant to the algorithm
-We now provide a formal definition for the same.
-Note that, as it is standard in the case of dynamic programming on tree decomposition, we will not compute all the partial solution at node $t$ but rather just store unique signatures of all the partial solutions at $t$.
-The signature $S$ of a partial solution $P=(\sigma,\tau,\mathcal{R},\mathcal{A})$ is $(\sigma|_{X_{t}},\tau|_{X_{t}},\mathcal{R},\mathcal{A})$.
-States of the Dynamic Programming Algorithms
-Definition 3.11 (Signature).
-A state of dynamic programming table will be a tuple $S=(\sigma,\tau,\mathcal{R},\mathcal{A})$ such that
-  • $\sigma:X_{t}\to X_{t}\cup\{\star\}$ (which will emulate a parent function),
-  • $\tau:X_{t}\to V(H_{t})\cup\{\star\}$ (which will emulate an originator function),
-  • $\mathcal{R}\subseteq R\times R\times\{\textsf{inv},\textsf{sat},\textsf{unsat}\}$ is a set of labeled pairs representing a label $\ell\in\{\textsf{inv},\textsf{sat},\textsf{unsat}\}$ assigned to the pair $(u,v)$, as per Definition 3.8, and
-  • $\mathcal{A}$ is the collection of pair of vertices in $X_{t}$ that are pseudo-adjacent with each other.
-The condition on functions ensures that at present $\sigma$ and $\tau$ are valid candidate to be a parent function and an originator function.
-For a node $t$ in tree decomposition $T$, we construct a dynamic-programming algorithm that sets the value $\textbf{d}[S,t]$ to True if and only if $S$ is a signature of some partial solution $P$ at node $t$.
-Otherwise, we set it to False.
-Initially, all entries are set to False.
-For any signature $S=(\sigma,\tau,\mathcal{R},\mathcal{A})$, if $\mathcal{R}$ contains an entry of the type $(u,v,\textsf{inv})$ then $\textbf{d}[S,t]$ is set to False and it will never be changed.
-3.1.3 Updating Dynamic Programming Tables
-We now proceed to describe a recursive algorithm used to update entries in the dynamic programming table for each node type in a nice tree-decomposition.
+Assume that $H$ is a contraction of $G$, and let $\mathcal{W}=\{W_{1},\dots,W_{|V(H)|}\}$ denote a hypothetical but fixed witness structure for this contraction.
+Recall that we refer to the unique vertex $h\in W_{i}\cap V(H)$ as the representative vertex of the witness set $W_{i}$.
+Definition 3.3 (Active Vertices).
+Let $t\in V(\mathcal{T})$ be a node with bag $X_{t}$.
+A representative vertex $h\in V(H)$ is active at $t$ if its corresponding witness set $W_{h}\in\mathcal{W}$ intersects with $X_{t}$, i.e., $W_{h}\cap X_{t}\neq\emptyset$.
+The set $A_{t}\subseteq V(H)$ is the collection of all active vertices at $t$.
+Observation 3.1.
+The number of active vertices $|A_{t}|$ is upper bounded by the size of the bag, $|A_{t}|\leq|X_{t}|$, as each active vertex corresponds to a unique witness set that intersects $X_{t}$.
+Furthermore, due to the connectivity properties of the witness sets, a representative vertex that becomes inactive in a bag $X_{t}$ (i.e., its witness set $W_{h}$ no longer intersects $X_{t}$) remains inactive in all subsequent bags in the traversal towards the root.
+Definition 3.4 (Witness Impression).
+For a given witness structure $\mathcal{W}$, the witness impression on the bag $X_{t}$ of node $t$ is the pair $(\mathcal{P},\tau)$, where:
+  • $\mathcal{P}$ is the partition of $X_{t}$ induced by the witness sets $\mathcal{W}$:
+    $\mathcal{P}=\{W\cap X_{t}\mid W\in\mathcal{W}\text{ and }W\cap X_{t}\neq\emptyset\}.$
+  • $\tau:\mathcal{P}\to(X_{t}\cap V(H))\cup\{\uparrow,\downarrow\}$ is a function that maps each part $P\in\mathcal{P}$ to the status of its unique representative vertex $h\in V(H)$:
+    $\tau(P)=\begin{cases}h&\text{if }h\in X_{t}\quad\text{(Representative in Bag)}\\ \downarrow&\text{if }h\in V(H_{t})\setminus X_{t}\quad\text{(Representative Forgotten)}\\ \uparrow&\text{if }h\in V(H)\setminus V(H_{t})\quad\text{(Representative Above)}\end{cases}$
+Definition 3.5 (Adjacency Sets).
+Consider a witness impression $(\mathcal{P},\tau)$ at node $t$.
+For any pair of distinct partitions $(P_{i},P_{j})\in\binom{\mathcal{P}}{2}$, let $h_{i}$ and $h_{j}$ be their respective representative vertices.
+We define the following two sets:
+  • $\mathcal{A}_{t}$ is the set of pseudo-adjacent pairs $(P_{i},P_{j})$ such that there is an edge in $G$ between $W_{i}\cap V(G_{t})$ and $W_{j}\cap V(G_{t})$, where $W_{i}\cap X_{t}=P_{i}$ and $W_{j}\cap X_{t}=P_{j}$.
+  • $\mathcal{R}_{t}$ is the set of anti-adjacent pairs $(P_{i},P_{j})$ such that the representatives $h_{i}$ and $h_{j}$ are non-adjacent in $H$, i.e., $(h_{i},h_{j})\notin E(H)$.
+In other words, $\mathcal{A}_{t}$ stores the pair of partitions such that the adjacency between the corresponding witness sets $W_{i}$ and $W_{j}$ is witnessed by some edge in $E(G_{t})$.
+It might be possible that both endpoints are completely in $V(G_{t})\setminus X_{t}$.
+Consider an example of witness sets $W_{1}$ and $W_{3}$ in Figure 3.
+Similarly, $\mathcal{R}_{t}$ stores the pair of partitions that should remain non-adjacent.
+This information is obvious if active vertices of the corresponding witness sets are in the bag.
+However, if one of both of them have been forgotten, the pairs in $\mathcal{R}_{t}$ provides the critical information.
+For example, in Figure 3, $\mathcal{R}_{t}$ contains a pair $(W_{1},W_{2})$ as their active vertices are non-adjacent in $H$.
+We are now in a state to describe our dynamic programming algorithm.
+The dynamic programming algorithm maintains a table $\texttt{dp}[t;S]$, where the signature $S$ at node $t$ is the tuple $(\mathcal{P},\tau,\mathcal{A}_{t},\mathcal{R}_{t})$.
+We set $\texttt{dp}[t;S]=\texttt{True}$ if and only if there exists an $H$-witness structure $\mathcal{W}$ of $G$ such that $S$ is the witness impression and adjacency sets derived from $\mathcal{W}$ at node $t$.
+Definition 3.6 (Valid Signature).
+Consider a partition $\mathcal{P}$ of $X_{t}$, a function $\tau:\mathcal{P}\mapsto(X_{t}\cap V(H))\cup\{\uparrow,\downarrow\}$, and two mutually disjoint sets $\mathcal{A}_{t},\mathcal{R}_{t}\subseteq\binom{\mathcal{P}}{2}$.
+The tuple $S=(\mathcal{P},\tau,\mathcal{A}_{t},\mathcal{R}_{t})$ is an valid signature if all of the following conditions hold:
+  1. 1.
+    (Cardinality) There is no partition $P_{i}\in\mathcal{P}$ that contains more than one vertex from $V(H)$.
+  2. 2.
+    (Unique Mapping) There do not exist two distinct partitions $P_{i},P_{j}\in\mathcal{P}$ such that $\tau(P_{i})=\tau(P_{j})\not\in\{\uparrow,\downarrow\}$.
+  3. 3.
+    (Adjacency Condition) There do not exist two partitions $P_{i},P_{j}$ such that $(P_{i},P_{j})\in\mathcal{R}_{t}$ and $P_{i}$ and $P_{j}$ are adjacent in $G$ (i.e., there is an edge between $P_{i}$ and $P_{j}$).
+  4. 4.
+    (Internal DP Constraint) There do not exist two partitions $P_{i},P_{j}$ such that $\tau(P_{i})=\uparrow$, $\tau(P_{j})=\downarrow$, and $P_{i}$ and $P_{j}$ are adjacent in $G$.
+An invalid signature is defined as a signature that does not satisfy the conditions of a valid signature.
+Lemma 3.2.
+For an invalid signature $S$, $\texttt{dp}[t;S]=\texttt{False}$.
+Proof.
+We demonstrate that if a signature $S=(\mathcal{P},\tau,\mathcal{A}_{t},\mathcal{R}_{t})$ satisfies any of the invalidity conditions, it cannot be derived from any valid $H$-witness structure $\mathcal{W}$ of $G$.
+Thus, the dp entry must be False.
+Let $\mathcal{W}$ be an arbitrary $H$-witness structure.
+  1. 1.
+    (Cardinality Error) If $P_{i}\in\mathcal{P}$ contains distinct vertices $h_{a},h_{b}\in V(H)$, then the witness set $W_{i}$ (of which $P_{i}$ is a fragment) contains multiple representative vertices.
+This violates the fundamental property of an $H$-witness structure, which requires each $W_{i}\in\mathcal{W}$ to contract to a single representative in $H$.
+  2. 2.
+    (Unique Mapping Error) If distinct partitions $P_{i},P_{j}\in\mathcal{P}$ map to the same representative $\tau(P_{i})=\tau(P_{j})=h\in X_{t}\cap V(H)$, then both $P_{i}$ and $P_{j}$ must belong to the same witness set $W_{h}$.
+Since $\mathcal{P}$ is the partition of $X_{t}$ induced by the distinct sets in $\mathcal{W}$, $P_{i}$ and $P_{j}$ cannot be distinct partitions, a contradiction.
+  3. 3.
+    (Adjacency Contradiction) If $(P_{i},P_{j})\in\mathcal{R}_{t}$, then their representatives $h_{i}$ and $h_{j}$ are non-adjacent in $H$.
+If, however, $P_{i}$ and $P_{j}$ are adjacent in $G$, the graph contraction rule requires $h_{i}$ and $h_{j}$ to be adjacent in $H$ to cover this edge.
+This direct contradiction implies no valid contraction exists for this signature.
+  4. 4.
+    (Internal DP Constraint) If partitions $P_{i}$ ($\tau(P_{i})=\uparrow$) and $P_{j}$ ($\tau(P_{j})=\downarrow$) are adjacent in $G$, the adjacency is between a witness set whose representative $h_{i}$ is yet to be introduced into the subgraph ($h_{i}\in V(H)\setminus V(H_{t})$) and a set whose representative $h_{j}$ has already been forgotten ($h_{j}\in V(H_{t})\setminus X_{t}$).
+This configuration represents the case when two witness sets are adjacent in $G$ but their representative vertices are not adjacent in $H$.
+Note that this situation can not be changed with additional vertices or edges.
+This structural inconsistency rules out a valid DP state.
+Therefore, no valid $H$-witness structure can produce an invalid signature $S$. ∎
+In rest of the section, whenever we mention a signature, we assume it to be valid according to Definition 3.6. We now present how the dynamic programming table is updated at each type of nodes.
 Leaf Node
-Let $t$ be a leaf node.
-Then $X_{t}=\emptyset$ and the valid partial solution $P$ is empty functions $\sigma,\tau,\mathcal{R}$ since their domain is a subset of $X_{t}$ which is trivially empty.
-Hence, the only valid signature consists of four empty sets.
+The following lemma is simple to prove.
 Lemma 3.3.
-If $t$ is a leaf node then set $\textbf{d}[S,t]=\texttt{True}$ for $S=(\emptyset,\emptyset,\emptyset,\emptyset)$, and $\textbf{d}[S^{\prime},t]=\texttt{False}$ for every other tuple $S^{\prime}$.
+Let $t$ be a leaf node. The DP entry is $\texttt{dp}[t;S]=\texttt{True}$ if and only if the signature is the empty tuple, $S=(\emptyset,\emptyset,\emptyset,\emptyset)$.
+Proof.
+The bag is $X_{t}=\emptyset$.
+Since the partition $\mathcal{P}$ is defined on $X_{t}$, $\mathcal{P}$ must be empty.
+Consequently, $\tau$ (with domain $\mathcal{P}$) and the adjacency sets $\mathcal{A}_{t}$ and $\mathcal{R}_{t}$ (defined on pairs of $\mathcal{P}$) must all be empty sets.
+As $G_{t}$ contains no vertices, the only signature consistent with any partial witness structure is the empty one.
+∎
 Introduce Vertex Node
 Let $t$ be an introduce vertex node in the tree decomposition $T$, with child $t^{\prime}$, such that $X_{t}=X_{t^{\prime}}\cup\{x\}$ for some vertex $x\in V(G)\setminus V(G_{t^{\prime}})$.
-Let $S=(\sigma,\tau,\mathcal{R},\mathcal{A})$ be a signature at node $t$, and let $S^{\prime}=(\sigma^{\prime},\tau^{\prime},\mathcal{R}^{\prime},\mathcal{A}^{\prime})$ be a signature at node $t^{\prime}$.
-Definition 3.12.
-We say that $S$ and $S^{\prime}$ are introduce-vertex compatible if the following conditions hold:
-  • ($C1$)
-    For every $y\in X_{t}\setminus\{x\}$, we have $\sigma(y)=\sigma^{\prime}(y)$ and $\tau(y)=\tau^{\prime}(y)$,
-  • ($C2$)
-    $\mathcal{R}^{\prime}\subseteq\mathcal{R}$ and $\mathcal{A}^{\prime}=\mathcal{A}$.
-Let $I(S)$ denote the set of all signatures at $t^{\prime}$ that are introduce-vertex compatible with $S$.
+Note that as we treat introduce edge node separately, at this stage, we do not have to consider any new edge.
+Let $S=(\mathcal{P},\tau,\mathcal{A}_{t},\mathcal{R}_{t})$ be a signature at node $t$, and let $S^{\prime}=(\mathcal{P}^{\prime},\tau^{\prime},\mathcal{A}_{t^{\prime}},\mathcal{R}_{t^{\prime}})$ be a signature at node $t^{\prime}$.
+Definition 3.7 (Introduce-Vertex Compatibility).
+Valid signatures $S=(\mathcal{P},\tau,\mathcal{A}_{t},\mathcal{R}_{t})$ at node $t$ and $S^{\prime}=(\mathcal{P}^{\prime},\tau^{\prime},\mathcal{A}_{t^{\prime}},\mathcal{R}_{t^{\prime}})$ at its child $t^{\prime}$ are introduce-vertex compatible if there exists a mapping $\phi:\mathcal{P}^{\prime}\to\mathcal{P}$ satisfying one of the following two cases:
+  1. (i)
+    $x$ joins an existing witness set: There exists a unique $P^{\prime}_{i}\in\mathcal{P}^{\prime}$ such that $\phi(P^{\prime}_{i})=P^{\prime}_{i}\cup\{x\}$, and for all other $P^{\prime}_{j}\in\mathcal{P}^{\prime}$, $\phi(P^{\prime}_{j})=P^{\prime}_{j}$.
+    • (C1)
+      If $x\in V(H)$, then $\tau(\phi(P^{\prime}_{i}))=x$ and $\tau^{\prime}(P^{\prime}_{i})=\uparrow$.
+Otherwise, $\tau(\phi(P^{\prime}_{j}))=\tau^{\prime}(P^{\prime}_{j})$ for all $P^{\prime}_{j}\in\mathcal{P}^{\prime}$.
+    • (C2)
+      $\mathcal{A}_{t}=\{(\phi(P^{\prime}_{a}),\phi(P^{\prime}_{b}))\mid(P^{\prime}_{a},P^{\prime}_{b})\in\mathcal{A}_{t^{\prime}}\}$.
+    • (C3)
+      $\mathcal{R}_{t}=\{(\phi(P^{\prime}_{a}),\phi(P^{\prime}_{b}))\mid(P^{\prime}_{a},P^{\prime}_{b})\in\mathcal{R}_{t^{\prime}}\}$.
+  2. (ii)
+    $x$ starts a new witness set: $\mathcal{P}=\mathcal{P}^{\prime}\cup\{P_{x}\}$ where $P_{x}=\{x\}$.
+The mapping $\phi$ is the identity $\phi(P^{\prime})=P^{\prime}$ for all $P^{\prime}\in\mathcal{P}^{\prime}$.
+    • (C1)
+      If $x\in V(H)$, then $\tau(P_{x})=x$. Otherwise, $\tau(P_{x})=\uparrow$.
+    • (C2)
+      $\mathcal{A}_{t}=\mathcal{A}_{t^{\prime}}$. (Note that $P_{x}$ cannot be part of any pseudo-adjacency as no edges are introduced).
+    • (C3)
+      $\mathcal{R}_{t}=\mathcal{R}_{t^{\prime}}\cup\{(P_{x},P_{j})\mid P_{j}\in\mathcal{P}^{\prime}\text{ and }\tau(P_{x}),\tau(P_{j})\in V(H)\text{ and }(\tau(P_{x}),\tau(P_{j}))\notin E(H)\}$.
 Lemma 3.4.
-Let $t$ be an introduce vertex node with child $t^{\prime}$, and let $X_{t}=X_{t^{\prime}}\cup\{x\}$ and $x\notin V(H)$.
-Let $S=(\sigma,\tau,\mathcal{R},\mathcal{A})$ be a signature at node $t$.
-Assign $\mathbf{d}[S,t]=\texttt{True}$ if and only if the following conditions are satisfied:
-  1. (i)
-    $\sigma(x)=\star$ and $\tau(x)=\star$,
-  2. (ii)
-    there exists a signature $S^{\prime}=(\sigma^{\prime},\tau^{\prime},\mathcal{R}^{\prime},\mathcal{A}^{\prime})\in I(S)$ such that $\mathbf{d}[S^{\prime},t^{\prime}]=\texttt{True}$,
-  3. (iii)
-    $\mathcal{R}=\mathcal{R}^{\prime}$ and $\mathcal{A}=\mathcal{A}^{\prime}$, i.e., no new elements are introduced in either of the sets.
+Let $t$ be an introduce vertex node with child $t^{\prime}$.
+For a valid signature $S$, $\texttt{dp}[t;S]=\texttt{True}$ if and only if there exists a signature $S^{\prime}$ such that $\texttt{dp}[t^{\prime};S^{\prime}]=\texttt{True}$ and $S,S^{\prime}$ are introduce-vertex compatible.
 Proof.
-$(\Leftarrow)$ Assume the three conditions of the lemma are met.
-By condition (ii), there exists a valid signature $S^{\prime}=(\sigma^{\prime},\tau^{\prime},\mathcal{R}^{\prime},\mathcal{A}^{\prime})$ at the child node $t^{\prime}$, which means there is a valid partial solution $P^{\prime}=(\sigma^{\prime}_{p},\tau^{\prime}_{p},\mathcal{R}^{\prime}_{p},\mathcal{A}^{\prime}_{p})$ for the subgraph $G_{t^{\prime}}$ that corresponds to $S^{\prime}$.
-We construct a new partial solution $P=(\sigma_{p},\tau_{p},\mathcal{R}_{p},\mathcal{A}_{p})$ for the subgraph $G_{t}$ by extending $P^{\prime}$ to include vertex $x$.
-For all $u\in V(G_{t^{\prime}})$, we set $\sigma_{p}(u)=\sigma^{\prime}_{p}(u)$ and $\tau_{p}(u)=\tau^{\prime}_{p}(u)$.
-For the new vertex $x\notin V(H)$, its parent must lie outside the current subtree.
-Thus, we set $\sigma_{p}(x)=\star$ and $\tau_{p}(x)=\star$, consistent with condition (i).
-Since no new edges are introduced involving $x$, we set $\mathcal{R}_{p}=\mathcal{R}^{\prime}_{p}$ and $\mathcal{A}_{p}=\mathcal{A}^{\prime}_{p}$, consistent with condition (iii).
-This constructed tuple $P$ is a valid partial solution for $G_{t}$, and its signature is precisely $S$.
-Therefore, $\mathbf{d}[S,t]$ is set to True.
-$(\Rightarrow)$ Now, assume that $\mathbf{d}[S,t]=\text{true}$.
-This means there exists a valid partial solution $P=(\sigma_{p},\tau_{p},\mathcal{R}_{p},\mathcal{A}_{p})$ for $G_{t}$ whose signature is $S$.
-We can define a partial solution $P^{\prime}$ for the smaller subgraph $G_{t^{\prime}}$ by simply restricting $P$.
-For all $u\in V(G_{t^{\prime}})$, we set $\sigma^{\prime}_{p}(u)=\sigma_{p}(u)$, $\tau^{\prime}_{p}(u)=\tau_{p}(u)$, $\mathcal{R}^{\prime}_{p}=\mathcal{R}_{p}$, and $\mathcal{A}^{\prime}_{p}=\mathcal{A}_{p}$.
-This new tuple $P^{\prime}$ is a valid partial solution for $G_{t^{\prime}}$, and its signature $S^{\prime}$ must exist in the DP table, meaning $\mathbf{d}[S^{\prime},t^{\prime}]=\text{true}$.
-This proves condition (ii).
-Since $x\notin V(H)$ and is a new, isolated vertex, it cannot be a parent to any other vertex in the subtree.
-This requires that in any valid partial solution, $\sigma_{p}(x)=\star$ and $\tau_{p}(x)=\star$, which proves condition (i).
-Finally, because no new edges or relationships were formed, the sets of requester pairs and pseudo-adjacencies remain unchanged from $t^{\prime}$ to $t$.
-Thus, $\mathcal{R}=\mathcal{R}^{\prime}$ and $\mathcal{A}=\mathcal{A}^{\prime}$, which proves condition (iii).
-∎
-Lemma 3.4.
-Let $t$ be an introduce vertex node with child $t^{\prime}$, and let $X_{t}=X_{t^{\prime}}\cup\{x\}$ and $x\in V(H)$.
-Let $S=(\sigma,\tau,\mathcal{R},\mathcal{A})$ be a signature at node $t$.
-Assign the value of $\mathbf{d}[S,t]=\texttt{True}$ if and only if the following conditions are met:
-  1. (i)
-    $\sigma(x)=x$ and $\tau(x)=x$,
-  2. (ii)
-    there exists a signature $S^{\prime}=(\sigma^{\prime},\tau^{\prime},\mathcal{R}^{\prime},\mathcal{A}^{\prime})\in I(S)$ such that $\mathbf{d}[S^{\prime},t^{\prime}]=\texttt{True}$,
-  3. (iii)
-    $\mathcal{A}=\mathcal{A}^{\prime}$, and every element in $\mathcal{R}\setminus\mathcal{R}^{\prime}$ is of the form
-    • $(x,u,\textsf{unsat})$ for some $u\in X_{t}$ such that $x$ and $\tau(u)$ are adjacent in $H$, or
-    • $(x,u,\textsf{sat})$ for some $u\in X_{t}$ such that $x$ and $\tau(u)$ are not adjacent in $H$.
-    The above condition implies that the only tuple introduced in $\mathcal{R}$ contains $x$ in it.
-Proof.
-$(\Leftarrow)$ Assume the three conditions of the lemma are met.
-By condition (ii), there exists a valid signature $S^{\prime}=(\sigma^{\prime},\tau^{\prime},\mathcal{R}^{\prime},\mathcal{A}^{\prime})$ at the child node $t^{\prime}$, which means there is a valid partial solution $P^{\prime}=(\sigma^{\prime}_{p},\tau^{\prime}_{p},\mathcal{R}^{\prime}_{p},\mathcal{A}^{\prime}_{p})$ for the subgraph $G_{t^{\prime}}$.
-We construct a new partial solution $P=(\sigma_{p},\tau_{p},\mathcal{R}_{p},\mathcal{A}_{p})$ for the subgraph $G_{t}$ by extending $P^{\prime}$ to include vertex $x$.
-For all $u\in V(G_{t^{\prime}})$, we set $\sigma_{p}(u)=\sigma^{\prime}_{p}(u)$ and $\tau_{p}(u)=\tau^{\prime}_{p}(u)$.
-For the new vertex $x\in V(H)$, it must be the root of its own witness set.
-Thus, we set $\sigma_{p}(x)=x$ and $\tau_{p}(x)=x$, which is consistent with condition (i).
-The pseudo-adjacency set is unchanged as no new edges from $G$ are introduced, so we set $\mathcal{A}_{p}=\mathcal{A}^{\prime}_{p}$, consistent with condition (iii).
-As $x$ is now a requester vertex, its relationships with other requesters in the bag must be recorded.
-Based on condition (iii), we define $\mathcal{R}_{p}=\mathcal{R}^{\prime}_{p}\cup\{(x,u,\textsf{unsat})\mid(x,\tau(u))\in E(H)\}\cup\{(x,u,\textsf{sat})\mid(x,\tau(u))\notin E(H)\}$.
-The constructed tuple $P$ is a valid partial solution for $G_{t}$, and its signature is precisely $S$.
-Therefore, $\mathbf{d}[S,t]$ is set to True.
-$(\Rightarrow)$ Now, assume that $\mathbf{d}[S,t]=\text{true}$.
-This means there exists a valid partial solution $P=(\sigma_{p},\tau_{p},\mathcal{R}_{p},\mathcal{A}_{p})$ for $G_{t}$ whose signature is $S$.
-We can define a partial solution $P^{\prime}$ for the smaller subgraph $G_{t^{\prime}}$ by simply restricting $P$.
-For all $u\in V(G_{t^{\prime}})$, we set $\sigma^{\prime}_{p}(u)=\sigma_{p}(u)$ and $\tau^{\prime}_{p}(u)=\tau_{p}(u)$.
-The sets $\mathcal{R}^{\prime}_{p}$ and $\mathcal{A}^{\prime}_{p}$ are simply the projections of $\mathcal{R}_{p}$ and $\mathcal{A}_{p}$ onto pairs of vertices in $X_{t^{\prime}}$.
-The resulting tuple $P^{\prime}$ is a valid partial solution for $G_{t^{\prime}}$ since it omits the new vertex $x$.
-Its signature $S^{\prime}$ must exist in the DP table, meaning $\mathbf{d}[S^{\prime},t^{\prime}]=\text{true}$.
-This proves condition (ii).
-Since $x\in V(H)$, by the properties of a valid witness structure, $x$ must be the root of its witness tree and its own originator.
-This means $\sigma(x)=x$ and $\tau(x)=x$, which proves condition (i).
-Finally, as no edges from $G$ are introduced, the pseudo-adjacency set for pairs in $X_{t}$ remains unchanged from $t^{\prime}$ to $t$.
-Any new entries in the requester set $\mathcal{R}$ must involve the new vertex $x$ and must be correctly classified based on its relationships to other requesters, as captured in condition (iii).
-∎
-Forget Vertex Node.
-Let $t$ be a forget vertex node in the tree decomposition $T$, with child $t^{\prime}$, such that $X_{t}=X_{t^{\prime}}\setminus\{x\}$ for some vertex $x\in X_{t^{\prime}}$.
-Let $S=(\sigma,\tau,\mathcal{R},\mathcal{A})$ be a signature at node $t$, and let $S^{\prime}=(\sigma^{\prime},\tau^{\prime},\mathcal{R}^{\prime},\mathcal{A}^{\prime})$ be a signature at node $t^{\prime}$.
-Definition 3.13.
-We say that $S$ and $S^{\prime}$ are forget-vertex compatible if the following conditions hold:
-  • (C1) The vertex $x$ is not an explorer with respect to $\sigma^{\prime}$, as per Definition 3.7.
-  • (C2) For every $y\in X_{t^{\prime}}\setminus\{x\}$, we have $\sigma(y)=\sigma^{\prime}(y)$ and $\tau(y)=\tau^{\prime}(y)$.
-  • (C3) $\mathcal{R}$ is the restriction of $\mathcal{R}^{\prime}$ to pairs not involving $x$, and $\mathcal{A}$ is the restriction of $\mathcal{A}^{\prime}$ to pairs not involving $x$.
-Let $F(S)$ denote the set of all signatures at $t^{\prime}$ that are forget-vertex compatible with $S$.
-Lemma 3.5.
-Let $t$ be a forget vertex node with child $t^{\prime}$, and let $X_{t}=X_{t^{\prime}}\setminus\{x\}$.
-Let $S=(\sigma,\tau,\mathcal{R},\mathcal{A})$ be a signature at node $t$.
-Then, $\mathbf{d}[S,t]=\texttt{True}$ if and only if the following conditions hold:
-  1. (i)
-    There exists a signature $S^{\prime}=(\sigma^{\prime},\tau^{\prime},\mathcal{R}^{\prime},\mathcal{A}^{\prime})\in F(S)$ such that $\mathbf{d}[S^{\prime},t^{\prime}]=\texttt{True}$.
-  2. (ii)
-    The requester set $\mathcal{R}^{\prime}$ contains no entry of the form $(x,u,\textsf{unsat})$ for any $u\in X_{t^{\prime}}\setminus\{x\}$.
-Proof.
-$(\Leftarrow)$ Assume the two conditions of the lemma are met.
-By condition (i), there exists a valid signature $S^{\prime}$ at $t^{\prime}$ corresponding to a valid partial solution $P^{\prime}$ for the subgraph $G_{t^{\prime}}$.
-We construct a new partial solution $P$ for $G_{t}$ by restricting $P^{\prime}$ to $V(G)\setminus\{x\}$.
-Since $P^{\prime}$ is valid, all adjacency requirements imposed on $G_{t^{\prime}}$ must hold.
-Condition (ii) explicitly guarantees that the vertex $x$ does not have any unsat adjacency requirements with any other vertex $u\in X_{t}$.
-Because $x$ is not an explorer (C1), all required parent/originator relations for $x$ must be contained within $G_{t^{\prime}}$.
-By C3, the signature $S$ is precisely the valid signature $S^{\prime}$ with all information about $x$ correctly removed.
-Since all constraints involving $x$ are resolved within $G_{t^{\prime}}$ and $P^{\prime}$ is valid, the restriction $P$ is a valid partial solution for $G_{t}$.
-Therefore, $\mathbf{d}[S,t]$ is set to True.
-$(\Rightarrow)$ Now, assume that $\mathbf{d}[S,t]=\text{true}$.
-This means there exists a valid partial solution $P$ for $G_{t}$ whose signature is $S$.
-This solution $P$ must be the restriction of some valid partial solution $P^{\prime}$ for $G_{t^{\prime}}$.
-This proves the existence of a compatible signature $S^{\prime}$ such that $\mathbf{d}[S^{\prime},t^{\prime}]=\text{true}$, proving condition (i).
-For $P^{\prime}$ to be a valid partial solution, all requirements within $G_{t^{\prime}}$ must be satisfied.
-If $\mathcal{R}^{\prime}$ contained an entry $(x,u,\textsf{unsat})$, the required edge in $H$ between $\tau(x)$ and $\tau(u)$ would not have been realized by the edges in $G_{t^{\prime}}$, contradicting the validity of $P^{\prime}$.
-Thus, no unsat entry involving $x$ can exist in $\mathcal{R}^{\prime}$.
-This proves condition (ii).
+$(\Rightarrow)$ If $\texttt{dp}[t;S]=\texttt{True}$ with witness structure $\mathcal{W}$, its restriction $\mathcal{W}^{\prime}$ to $G_{t^{\prime}}$ induces a signature $S^{\prime}$.
+Since $x$ has no incident edges in $G_{t}$, the adjacency set $\mathcal{A}_{t}$ is simply the projection of $\mathcal{A}_{t^{\prime}}$.
+If $x$ joins an existing part $P^{\prime}_{i}$ (Case i), the mapping $\phi$ updates the part.
+If $x$ starts a new part (Case ii), $\mathcal{R}_{t}$ is correctly initialized to reflect non-adjacency requirements in $H$.
+In both cases, $S$ and $S^{\prime}$ are compatible.
+$(\Leftarrow)$ Given a witness structure for $G_{t^{\prime}}$ inducing $S^{\prime}$, we extend it by adding $x$ to a witness set as specified by $\phi$.
+Since $x$ is not incident to any edges in $G_{t}$, no new pseudo-adjacencies are created, making the update $\mathcal{A}_{t}=\mathcal{A}_{t^{\prime}}$ consistent.
+The updates to $\tau$ and $\mathcal{R}_{t}$ ensure that representative identities and non-adjacency requirements are correctly tracked.
+Thus, the extended structure is a valid partial $H$-witness structure for $G_{t}$.
 ∎
 Introduce Edge Node
-Let $t$ be an introduce edge node with child $t^{\prime}$, so that $X_{t}=X_{t^{\prime}}$, and let the edge $(u,v)\in E(G)$ be introduced at $t$, with $u,v\in X_{t}$.
-Let $S=(\sigma,\tau,\mathcal{R},\mathcal{A})$ be a candidate signature at $t$.
-Definition 3.14.
-We say that $S$ is introduce-edge compatible with respect to its counterpart at $t^{\prime}$, denoted $S^{\prime}=(\sigma^{\prime},\tau^{\prime},\mathcal{R}^{\prime},\mathcal{A}^{\prime})$, if the following hold:
-  • (C1) For every $x\in X_{t}\setminus\{u,v\}$, we have $\sigma(x)=\sigma^{\prime}(x)$ and $\tau(x)=\tau^{\prime}(x)$.
-  • (C2) At most one of the following statements is true for the values of $\sigma$ and $\tau$ at $u$ and $v$:
-    • (C2.1) No change in values. Formally, $\sigma^{\prime}(u)=\sigma(u)$, $\tau^{\prime}(u)=\tau(u)$ and $\sigma^{\prime}(v)=\sigma(v)$, $\tau^{\prime}(v)=\tau(v)$.
-    • (C2.2) $v$’s set merges into $u$’s set.
-Formally, $\sigma(u)=\sigma^{\prime}(u)$, $\tau(u)=\tau^{\prime}(u)$, $\sigma(v)\neq\sigma^{\prime}(v)$, and $\tau(v)=\tau(u)$.
-This change represents $u$ becoming the parent of $v$ or the root of $v$’s witness tree.
-    • (C2.3) $u$’s set merges into $v$’s set. Formally, $\sigma(v)=\sigma^{\prime}(v)$, $\tau(v)=\tau^{\prime}(v)$, $\sigma(u)\neq\sigma^{\prime}(u)$, and $\tau(u)=\tau(v)$.
-  • (C3) $\mathcal{R}^{\prime}\setminus\mathcal{R}\subseteq\{(u,v,\textsf{unsat}),(v,u,\textsf{unsat})\}$ and $\mathcal{A}^{\prime}\subseteq\mathcal{A}$.
-Let $I_{e}(S)$ denote the set of all signatures at $t^{\prime}$ that are introduce-edge compatible with $S$.
+Let $t$ be an introduce edge node in the tree decomposition $T$, with child $t^{\prime}$, such that $X_{t}=X_{t^{\prime}}$, and an edge $e=(u,v)\in E(G)$ is introduced at $t$, where $u,v\in X_{t}$.
+Let $P_{u},P_{v}\in\mathcal{P}$ be the parts containing $u$ and $v$, respectively.
+Let $S=(\mathcal{P},\tau,\mathcal{A}_{t},\mathcal{R}_{t})$ be a signature at node $t$, and $S^{\prime}=(\mathcal{P}^{\prime},\tau^{\prime},\mathcal{A}_{t^{\prime}},\mathcal{R}_{t^{\prime}})$ be a signature at node $t^{\prime}$.
+Definition 3.8 (Introduce-Edge Compatibility).
+Valid signatures $S=(\mathcal{P},\tau,\mathcal{A}_{t},\mathcal{R}_{t})$ at node $t$ and $S^{\prime}=(\mathcal{P}^{\prime},\tau^{\prime},\mathcal{A}_{t^{\prime}},\mathcal{R}_{t^{\prime}})$ at child $t^{\prime}$ are introduce-edge compatible if all of the following conditions hold:
+  • (C1)
+    $\mathcal{P}=\mathcal{P}^{\prime}$, $\tau=\tau^{\prime}$, and $\mathcal{R}_{t}=\mathcal{R}_{t^{\prime}}$.
+  • (C2)
+    Adjacency Update:
+    1. (i)
+      If $P_{u}=P_{v}$, then $\mathcal{A}_{t}=\mathcal{A}_{t^{\prime}}$.
+    2. (ii)
+      If $P_{u}\neq P_{v}$, then $\mathcal{A}_{t}=\mathcal{A}_{t^{\prime}}\cup\{(P_{u},P_{v})\}$.
 Lemma 3.5.
-Let $t$ be an introduce edge node where edge $(u,v)$ is introduced, with $u,v\in V(H)$.
-A signature $S$ is valid at node $t$ if and only if there exists a signature $S^{\prime}\in I_{e}(S)$ such that $\mathbf{d}[S^{\prime},t^{\prime}]=\texttt{True}$, and the following conditions hold:
-  1. (i)
-    Condition C2.1 holds.
-  2. (ii)
-    The pseudo-adjacency set is unchanged: $\mathcal{A}=\mathcal{A}^{\prime}$.
-  3. (iii)
-    The requester set is updated only for $(u,v)$: $\mathcal{R}=\mathcal{R}^{\prime}\setminus\{(u,v,\textsf{unsat})\}\cup\{(u,v,\textsf{sat})\}$, if $(u,v,\textsf{unsat})\in\mathcal{R}^{\prime}$.
-Otherwise $\mathcal{R}=\mathcal{R}^{\prime}$.
+Let $t$ be an introduce edge node with child $t^{\prime}$, and let $e=(u,v)$ be the introduced edge.
+Then, $\texttt{dp}[t;S]=\texttt{True}$ if and only if there exists a signature $S^{\prime}\in\mathcal{S}_{t^{\prime}}$ such that $\texttt{dp}[t^{\prime};S^{\prime}]=\texttt{True}$ and $S$ and $S^{\prime}$ are introduce-edge compatible.
 Proof.
-($\Leftarrow$) Assume the conditions are met.
-Condition (i) must hold since $u,v\in V(H)$, meaning they must be the roots of their witness sets, which cannot merge due to a $G$-edge.
-By the premise, a valid partial solution $P^{\prime}$ exists for $G_{t^{\prime}}$.
-We construct $P$ for $G_{t}$ by adding $(u,v)$.
-Since the functions $\sigma_{p},\tau_{p}$ and $\mathcal{A}_{p}$ are unchanged (condition (i) and (ii)), the only effect is to satisfy the $H$-requirement between $\tau(u)$ and $\tau(v)$.
-Condition (iii) correctly formalizes this satisfaction, updating an unsat requirement to sat in $\mathcal{R}$.
-Thus $P$ is a valid partial solution for $G_{t}$.
-($\Rightarrow$) If $\mathbf{d}[S,t]=\text{true}$, a valid $P$ exists, derived from a valid $P^{\prime}$ at $t^{\prime}$.
-Since $u,v\in V(H)$, the edge $(u,v)$ cannot cause a witness set merge, so the functions must remain the same, proving condition (i).
-The edge is a $G$-edge, not a pseudo-adjacency, so $\mathcal{A}$ remains $\mathcal{A}^{\prime}$, proving condition (ii).
-Its only effect is on the $H$-contraction requirement, which means if $(u,v,\textsf{unsat})$ was in $\mathcal{R}^{\prime}$, it must be resolved to sat in $\mathcal{R}$.
-If it was already sat or inv, $\mathcal{R}$ must be the same as $\mathcal{R}^{\prime}$.
-This verifies condition (iii).
+$(\Rightarrow)$ If $\texttt{dp}[t;S]=\texttt{True}$ with witness structure $\mathcal{W}$, then $\mathcal{W}$ restricted to $G_{t^{\prime}}$ induces a signature $S^{\prime}$.
+Since $X_{t}=X_{t^{\prime}}$, only the adjacency set $\mathcal{A}_{t}$ can change.
+If the introduced edge $e=\{u,v\}$ connects different parts $P_{u}$ and $P_{v}$, it witnesses a new pseudo-adjacency $(P_{u},P_{v})\in\mathcal{A}_{t}$, while all other components remain identical to $S^{\prime}$.
+$(\Leftarrow)$ Given a witness structure for $G_{t^{\prime}}$ inducing $S^{\prime}$, we add the edge $e=\{u,v\}$ to form $G_{t}$.
+If $P_{u}\neq P_{v}$, this edge creates a new adjacency between witness sets, which is correctly reflected in $\mathcal{A}_{t}$ by the compatibility condition.
+Because $S$ is valid, this new adjacency does not violate any non-adjacency constraints in $\mathcal{R}_{t}$, so the structure remains a valid partial $H$-witness structure for $G_{t}$.
 ∎
+Forget Vertex Node
+Let $t$ be a forget vertex node in the tree decomposition $T$, with child $t^{\prime}$, such that $X_{t}=X_{t^{\prime}}\setminus\{x\}$ for some vertex $x\in X_{t^{\prime}}$.
+Let $S=(\mathcal{P},\tau,\mathcal{A}_{t},\mathcal{R}_{t})$ be a signature at node $t$, and $S^{\prime}=(\mathcal{P}^{\prime},\tau^{\prime},\mathcal{A}_{t^{\prime}},\mathcal{R}_{t^{\prime}})$ be a signature at node $t^{\prime}$.
+Let $P_{x}\in\mathcal{P}^{\prime}$ be the part containing $x$, and $P_{x}^{*}=P_{x}\setminus\{x\}$.
+Definition 3.9 (Forget-Vertex Compatibility).
+Valid signatures $S=(\mathcal{P},\tau,\mathcal{A}_{t},\mathcal{R}_{t})$ at node $t$ and $S^{\prime}=(\mathcal{P}^{\prime},\tau^{\prime},\mathcal{A}_{t^{\prime}},\mathcal{R}_{t^{\prime}})$ at child $t^{\prime}$ are forget-vertex compatible if one of the following two cases holds:
+  • $x$ is not the only vertex in its part ($P_{x}\neq\{x\}$):
+    • (C1)
+      $\mathcal{P}=(\mathcal{P}^{\prime}\setminus\{P_{x}\})\cup\{P_{x}^{*}\}$.
+    • (C2)
+      For all $P\in\mathcal{P}^{\prime}\setminus\{P_{x}\}$, $\tau(P)=\tau^{\prime}(P)$. For the resulting part $P_{x}^{*}$:
+      1. 1.
+        If $\tau^{\prime}(P_{x})=x$, then $\tau(P_{x}^{*})=\downarrow$.
+      2. 2.
+        Otherwise, $\tau(P_{x}^{*})=\tau^{\prime}(P_{x})$.
+    • (C3)
+      $\mathcal{A}_{t}$ and $\mathcal{R}_{t}$ are the restrictions of $\mathcal{A}_{t^{\prime}}$ and $\mathcal{R}_{t^{\prime}}$ to $\mathcal{P}$.
+  • $x$ is the last vertex in its part ($P_{x}=\{x\}$):
+    • (C1)
+      $\mathcal{P}=\mathcal{P}^{\prime}\setminus\{P_{x}\}$.
+    • (C2)
+      For all $P\in\mathcal{P}$, $\tau(P)=\tau^{\prime}(P)$.
+    • (C3)
+      $\mathcal{A}_{t}$ and $\mathcal{R}_{t}$ are the restrictions of $\mathcal{A}_{t^{\prime}}$ and $\mathcal{R}_{t^{\prime}}$ to $\mathcal{P}$.
+    • (C4)
+      For all $P_{j}\in\mathcal{P}$, if $(P_{x},P_{j})\notin\mathcal{R}_{t^{\prime}}$, then $(P_{x},P_{j})\in\mathcal{A}_{t^{\prime}}$.
 Lemma 3.6.
-Let $t$ be an introduce edge node where edge $(u,v)$ is introduced, with $u\in V(H)$ and $v\notin V(H)$.
-A signature $S$ is valid at node $t$ if and only if there exists a signature $S^{\prime}\in I_{e}(S)$ such that $\mathbf{d}[S^{\prime},t^{\prime}]=\texttt{True}$, and one of the following holds:
-  1. (i)
-    Condition C2.1 holds and $\mathcal{R}=\mathcal{R}^{\prime}$ and $\mathcal{A}=\mathcal{A}^{\prime}\cup\{(u,v)\}$.
-  2. (ii)
-    Condition C2.2 holds and $\mathcal{R}=\mathcal{R}^{\prime}$ and $\mathcal{A}=\mathcal{A}^{\prime}$.
+Let $t$ be a forget vertex node with child $t^{\prime}$.
+For a valid signature $S$, $\texttt{dp}[t;S]=\texttt{True}$ if and only if there exists a signature $S^{\prime}$ such that $\texttt{dp}[t^{\prime};S^{\prime}]=\texttt{True}$ and $S,S^{\prime}$ are forget-vertex compatible.
 Proof.
-($\Leftarrow$) Assume one of the conditions holds. A valid $S^{\prime}$ exists by the premise.
-  1. 1.
-    If (i) holds, the functions are preserved (C2.1), meaning no witness set merge occurs.
-The edge $(u,v)$ is thus recorded as a new pseudo-adjacency in $\mathcal{A}$, forming a valid $P$ for $G_{t}$.
-  2. 2.
-    If (ii) holds, the functions change according to C2.2, meaning the witness set containing $v$ merges into $u$’s set.
-The edge $(u,v)$ is absorbed into the new witness structure and does not become a pseudo-adjacency, so $\mathcal{A}$ remains $\mathcal{A}^{\prime}$.
-This forms a valid $P$ for $G_{t}$.
-($\Rightarrow$) If $\mathbf{d}[S,t]=\text{true}$, a valid $P$ exists.
-Since $u\in V(H)$ and $v\notin V(H)$, $u$ must be a requester and $v$ cannot be the root of a witness set.
-The new edge $(u,v)$ must either form a pseudo-adjacency (case (i)) or cause $v$’s set to be absorbed by $u$’s (case (ii)).
-Case (i) requires function stability (C2.1).
-Case (ii) requires the witness set merge, which corresponds to the $\sigma$ and $\tau$ changes in C2.2.
-These two cases are exhaustive and mutually exclusive for a valid solution.
-∎
-Lemma 3.8.
-Let $t$ be an introduce edge node where edge $(u,v)$ is introduced, with $u\notin V(H)$ and $v\in V(H)$.
-A signature $S$ is valid at node $t$ if and only if there exists a signature $S^{\prime}\in I_{e}(S)$ such that $\mathbf{d}[S^{\prime},t^{\prime}]=\texttt{True}$, and one of the following holds:
-  1. (i)
-    Condition C2.1 holds and $\mathcal{A}=\mathcal{A}^{\prime}\cup\{(u,v)\}$.
-  2. (ii)
-    Condition C2.3 holds and $\mathcal{R}=\mathcal{R}^{\prime}$ and $\mathcal{A}=\mathcal{A}^{\prime}$.
-Proof.
-This is symmetric to Lemma 3.7, with $u$ and $v$ roles swapped.
-Case (i) is for pseudo-adjacency (no merge), and Case (ii) is for $u$’s set merging into $v$’s set (C2.3).
-The proof follows the same logic.
-∎
-Lemma 3.9.
-Let $t$ be an introduce edge node where edge $(u,v)$ is introduced, with $u\notin V(H)$ and $v\notin V(H)$.
-A signature $S$ is valid at node $t$ if and only if there exists a signature $S^{\prime}\in I_{e}(S)$ such that $\mathbf{d}[S^{\prime},t^{\prime}]=\texttt{True}$, and one of the following holds:
-  1. (i)
-    Condition C2.1 holds and $\mathcal{A}=\mathcal{A}^{\prime}\cup\{(u,v)\}$.
-  2. (ii)
-    Condition C2.2 holds and $\mathcal{R}=\mathcal{R}^{\prime}$ and $\mathcal{A}=\mathcal{A}^{\prime}$.
-  3. (iii)
-    Condition C2.3 holds and $\mathcal{R}=\mathcal{R}^{\prime}$ and $\mathcal{A}=\mathcal{A}^{\prime}$.
-Proof.
-This case combines the possibilities from the previous two lemmas.
-Since neither $u$ nor $v$ are $H$-vertices, the edge $(u,v)$ can either: (i) be a pseudo-adjacency (functions unchanged, C2.1), or (ii) cause $v$’s set to merge into $u$’s (C2.2), or (iii) cause $u$’s set to merge into $v$’s (C2.3).
-In the merge cases ((ii) and (iii)), the edge is absorbed into the witness structure, leaving $\mathcal{A}$ unchanged.
-The proof follows the same construction and existence logic as the previous cases.
+$(\Rightarrow)$ If $\texttt{dp}[t;S]=\texttt{True}$ with witness structure $\mathcal{W}$, then $\mathcal{W}$ induces a signature $S^{\prime}$ for $G_{t^{\prime}}$ containing $x$.
+Since $G_{t}=G_{t^{\prime}}$, the only difference is the restriction of the partition to $X_{t}\setminus\{x\}$.
+If $x$ was the last vertex of its part in the bag ($P_{x}=\{x\}$), any witness set $W_{j}$ required to be adjacent to $W_{x}$ in $H$ (where $(P_{x},P_{j})\notin\mathcal{R}_{t^{\prime}}$) must have already established an adjacency in $G_{t^{\prime}}$, satisfying (C4).
+$(\Leftarrow)$ A valid witness structure for $G_{t^{\prime}}$ remains valid for $G_{t}$.
+The compatibility cases ensure $S$ correctly tracks the representative status after $x$ is removed.
+If $P_{x}$ is removed from the bag (Case 2), condition (C4) guarantees that $W_{x}$ has already satisfied all required adjacencies with other witness sets.
+Since $S$ is valid, the forgotten witness set $W_{x}$ does not violate any global $H$-contraction requirements, so $\texttt{dp}[t;S]=\texttt{True}$.
 ∎
 Join Node
-Let $t$ be a join node in the tree decomposition $T$, with children $t_{1}$ and $t_{2}$, such that $X_{t}=X_{t_{1}}=X_{t_{2}}$.
-Let $S=(\sigma,\tau,\mathcal{R},\mathcal{A})$ be a signature at node $t$, and let $S_{1}=(\sigma_{1},\tau_{1},\mathcal{R}_{1},\mathcal{A}_{1})$ and $S_{2}=(\sigma_{2},\tau_{2},\mathcal{R}_{2},\mathcal{A}_{2})$ be signatures at nodes $t_{1}$ and $t_{2}$ respectively.
-Definition 3.15.
-We say that $S$, $S_{1}$, and $S_{2}$ are join-node compatible if the following conditions hold:
-  • (C1) For every $y\in X_{t}$, we have $\sigma(y)=\sigma_{1}(y)=\sigma_{2}(y)$ and $\tau(y)=\tau_{1}(y)=\tau_{2}(y)$.
-  • (C2) The signature sets are combined by union: $\mathcal{R}=\mathcal{R}_{1}\cup\mathcal{R}_{2}$ and $\mathcal{A}=\mathcal{A}_{1}\cup\mathcal{A}_{2}$.
-Let $J(S)$ denote the set of all pairs of signatures $(S_{1},S_{2})$ at $t_{1}$ and $t_{2}$ respectively, that satisfy (C1) and (C2) with $S$.
+Let $t$ be a join node with children $t_{L}$ and $t_{R}$, such that $X_{t}=X_{t_{L}}=X_{t_{R}}$.
+Definition 3.10 (Join-Node Compatibility).
+Valid signatures $S=(\mathcal{P},\tau,\mathcal{A}_{t},\mathcal{R}_{t})$ at node $t$, $S_{L}=(\mathcal{P}_{L},\tau_{L},\mathcal{A}_{L},\mathcal{R}_{L})$ and $S_{R}=(\mathcal{P}_{R},\tau_{R},\mathcal{A}_{R},\mathcal{R}_{R})$ at children $t_{L},t_{R}$ are join-compatible if the following conditions hold:
+  • (C1)
+    $\mathcal{P}=\mathcal{P}_{L}=\mathcal{P}_{R}$. Furthermore, for every $P\in\mathcal{P}$:
+    • If $\tau_{L}(P)\in V(H)$ and $\tau_{R}(P)=\uparrow$, then $\tau(P)=\tau_{L}(P)$.
+    • If $\tau_{R}(P)\in V(H)$ and $\tau_{L}(P)=\uparrow$, then $\tau(P)=\tau_{R}(P)$.
+    • If $\tau_{L}(P)=\downarrow$ or $\tau_{R}(P)=\downarrow$, then $\tau(P)=\downarrow$.
+    • Otherwise, $\tau_{L}(P)=\tau_{R}(P)=\tau(P)$.
+  • (C2)
+    $\mathcal{A}_{t}=\mathcal{A}_{L}\cup\mathcal{A}_{R}$.
+  • (C3)
+    $\mathcal{R}_{t}=\mathcal{R}_{L}=\mathcal{R}_{R}$.
 Lemma 3.7.
-Let $t$ be a join node with children $t_{1}$ and $t_{2}$, and let $X_{t}=X_{t_{1}}=X_{t_{2}}$.
-Let $S=(\sigma,\tau,\mathcal{R},\mathcal{A})$ be a signature at node $t$.
-Then, $\mathbf{d}[S,t]=\texttt{True}$ if and only if there exists a pair of signatures $(S_{1},S_{2})\in J(S)$, where $S_{1}=(\sigma_{1},\tau_{1},\mathcal{R}_{1},\mathcal{A}_{1})$ is a signature at $t_{1}$ and $S_{2}=(\sigma_{2},\tau_{2},\mathcal{R}_{2},\mathcal{A}_{2})$ is a signature at $t_{2}$, such that:
-  1. (i)
-    $\mathbf{d}[S_{1},t_{1}]=\texttt{True}$ and $\mathbf{d}[S_{2},t_{2}]=\texttt{True}$.
-  2. (ii)
-    The sets $\mathcal{A}$ and $\mathcal{R}$ are formed by combining the corresponding sets from $S_{1}$ and $S_{2}$ according to the following rules for every pair $(u,v)\in X_{t}\times X_{t}$:
-    • $\mathcal{A}=\mathcal{A}_{1}\cup\mathcal{A}_{2}$.
-    • $\mathcal{R}$ is the union of all combined requester entries, $\mathcal{R}=\bigcup_{u,v\in X_{t}}\mathcal{R}(u,v)$, where $\mathcal{R}(u,v)$ is determined by the $\oplus$ and $\otimes$ operators:
-      • Requester Merge ($\oplus$): For $(u,v,\ell_{1})\in\mathcal{R}_{1}$ and $(u,v,\ell_{2})\in\mathcal{R}_{2}$:
-[…]
-      • Pseudo-Adjacency Satisfies ($\otimes$): For $(u,v,\ell_{1})\in\mathcal{R}_{1}$ and $(u,v)\in\mathcal{A}_{2}$:
-        $\ell_{1}\otimes\mathcal{A}_{2}=\begin{cases}\textsf{sat}&\text{if }\ell_{1}=\textsf{unsat}\\ \ell_{1}&\text{if }\ell_{1}\in\{\textsf{sat},\textsf{inv}\}\end{cases}$
-        (A symmetric rule applies for $\mathcal{R}_{2}\otimes\mathcal{A}_{1}$.)
-      The resulting entry for $(u,v)$ in $\mathcal{R}$ is $\ell_{1}\oplus\ell_{2}$ (if both exist), or the result of the $\otimes$ operations, prioritizing sat over unsat and unsat over non-existence, and rejecting inv.
+Let $t$ be a join node with children $t_{L},t_{R}$.
+For a valid signature $S$, $\texttt{dp}[t;S]=\texttt{True}$ if and only if there exist signatures $S_{L},S_{R}$ such that $\texttt{dp}[t_{L};S_{L}]=\texttt{True}$, $\texttt{dp}[t_{R};S_{R}]=\texttt{True}$, and $S,S_{L},S_{R}$ are join-compatible.
 Proof.
-$(\Leftarrow)$ Assume conditions (i) and (ii) hold.
-By condition (i), there exist valid partial solutions $P_{1}=(\sigma_{p1},\tau_{p1},\mathcal{R}_{p1},\mathcal{A}_{p1})$ for $G_{t_{1}}$ and $P_{2}=(\sigma_{p2},\tau_{p2},\mathcal{R}_{p2},\mathcal{A}_{p2})$ for $G_{t_{2}}$.
-We construct a new partial solution $P=(\sigma_{p},\tau_{p},\mathcal{R}_{p},\mathcal{A}_{p})$ for $G_{t}=G_{t_{1}}\cup G_{t_{2}}$.
-The functions $\sigma_{p}$ and $\tau_{p}$ are defined piece-wise by the union of $\sigma_{p1},\tau_{p1}$ and $\sigma_{p2},\tau_{p2}$.
-Since $X_{t}=X_{t_{1}}=X_{t_{2}}$, condition (C1) ensures the functions agree on $X_{t}$, guaranteeing that the union is a valid structure for the witness sets.
-The set of edges $E(G_{t})=E(G_{t_{1}})\cup E(G_{t_{2}})$.
-  • Pseudo-Adjacency ($\mathcal{A}$): The pseudo-adjacencies in $P$ are exactly the union of those in $P_{1}$ and $P_{2}$, as $G_{t_{1}}$ and $G_{t_{2}}$ share no edges.
-Condition (ii) enforces $\mathcal{A}=\mathcal{A}_{1}\cup\mathcal{A}_{2}$.
-  • Requester Set ($\mathcal{R}$): The requester set tracks $H$-requirements between the originators of vertices in $X_{t}$.
-A requirement is satisfied in $G_{t}$ if it is satisfied in $G_{t_{1}}$ or $G_{t_{2}}$, or if the necessary edge is provided by a pseudo-adjacency in the other subgraph.
-    • The $\oplus$ operator correctly determines the status based on the union of $G_{t_{1}}$ and $G_{t_{2}}$: an unsat requirement becomes sat if the edge is found in the other subgraph (represented by an sat in the other $\mathcal{R}$ set).
-    • The $\otimes$ operator correctly models the satisfaction of an unsat requirement in one subgraph ($\mathcal{R}_{1}$ or $\mathcal{R}_{2}$) by a pseudo-adjacency recorded in the other subgraph ($\mathcal{A}_{2}$ or $\mathcal{A}_{1}$).
-    Condition (ii) correctly computes the final status for $\mathcal{R}$ ensuring all constraints are properly inherited and resolved.
-Since $P_{1}$ and $P_{2}$ are valid, and $P$ is constructed consistently with the union rules, $P$ is a valid partial solution for $G_{t}$.
-The signature of $P$ is $S$, thus $\mathbf{d}[S,t]=\texttt{True}$.
-$(\Rightarrow)$ Assume $\mathbf{d}[S,t]=\text{true}$.
-There exists a valid partial solution $P$ for $G_{t}$ whose signature is $S$.
-By the definition of $G_{t}$ and tree decomposition, $P$ must be the result of combining two valid partial solutions $P_{1}$ for $G_{t_{1}}$ and $P_{2}$ for $G_{t_{2}}$.
-  • Let $S_{1}$ and $S_{2}$ be the signatures of $P_{1}$ and $P_{2}$, respectively. Since $P_{1}$ and $P_{2}$ are valid, condition (i) must hold.
-  • Since $P$ is a valid union of $P_{1}$ and $P_{2}$ and $X_{t}$ is the intersection, the function values must agree on $X_{t}$, which verifies (C1) and ensures $(S_{1},S_{2})\in J(S)$.
-  • The pseudo-adjacency set $\mathcal{A}$ in $S$ must be the union of the pseudo-adjacencies recorded in $S_{1}$ and $S_{2}$ as the edges are disjoint, verifying $\mathcal{A}=\mathcal{A}_{1}\cup\mathcal{A}_{2}$.
-  • The final status of any $H$-requirement $(u,v)$ in $\mathcal{R}$ must be the result of combining the status from $S_{1}$ and $S_{2}$, accounting for satisfaction via: (1) internal edges in $G_{t_{1}}$, (2) internal edges in $G_{t_{2}}$, or (3) pseudo-adjacencies in the opposite subgraph.
-This requires the exact combination defined by the $\oplus$ and $\otimes$ operators in condition (ii) to hold.
-Since a valid $P$ exists, a pair $(S_{1},S_{2})$ satisfying all specified conditions must exist. ∎
-3.1.4 Putting it all together.
-We are now in a position to present the proof of Theorem 1 which states that Labeled Contraction admits an algorithm with running time $2^{\mathcal{O}(\textsf{tw}^{2})}\cdot|V(G)|^{\mathcal{O}(1)}$.
+$(\Rightarrow)$ If $\texttt{dp}[t;S]=\texttt{True}$, the witness structure $\mathcal{W}$ for $G_{t}$ restricted to $G_{t_{L}}$ and $G_{t_{R}}$ must induce $S_{L}$ and $S_{R}$.
+Since $V(G_{t_{L}})\cap V(G_{t_{R}})=X_{t}$, the bag partitions must be identical.
+The status $\tau(P)$ correctly merges the representative’s state (present, missing, or forgotten) from both branches.
+Since $E(G_{t})=E(G_{t_{L}})\cup E(G_{t_{R}})$, any set-to-set adjacency witnessed in $G_{t}$ must appear in $\mathcal{A}_{L}\cup\mathcal{A}_{R}$.
+$(\Leftarrow)$ Given compatible $S_{L},S_{R}$, we merge their witness structures by taking the union of corresponding witness sets.
+Compatibility ensures they agree on the bag partition and representatives.
+The union $\mathcal{A}_{t}=\mathcal{A}_{L}\cup\mathcal{A}_{R}$ accounts for all edges in $G_{t}$.
+Since $S$ is valid, this union satisfies all non-adjacency requirements in $\mathcal{R}_{t}$, making the merged structure a valid partial $H$-witness structure for $G_{t}$.
+∎
+We are finally at a position to present proof of algorithm mentioned in Theorem 1.
 Proof.
-The proof relies on the Dynamic Programming (DP) approach defined by the signatures $S=(\sigma,\tau,\mathcal{R},\mathcal{A})$ and the update laws (Lemmas 3.2 through 3.10).
-The overall running time is determined by the number of nodes in the tree decomposition, the size of the DP state space, and the time required for each transition.
-At any node $t$ of the tree decomposition, the size of the bag $|X_{t}|$ is at most $\textsf{tw}+1$.
-The number of distinct signatures $S=(\sigma,\tau,\mathcal{R},\mathcal{A})$ stored at node $t$ is bounded by considering each component: The Parent Function $\sigma:X_{t}\to X_{t}\cup\{\star\}$ and the Originator Function $\tau:X_{t}\to V(H)\cup\{\star\}$ each contribute $2^{\mathcal{O}(\textsf{tw}\log\textsf{tw})}$ states.
-For $\tau$, this is because only $\mathcal{O}(\textsf{tw})$ vertices in $V(H)$ are relevant to the bag $X_{t}$.
-The Requester Set $\mathcal{R}$ encodes a status from $\{\textsf{sat},\textsf{unsat},\textsf{inv}\}$ for $\mathcal{O}(\textsf{tw}^{2})$ pairs, resulting in $3^{\mathcal{O}(\textsf{tw}^{2})}=2^{\mathcal{O}(\textsf{tw}^{2})}$ choices.
-The Pseudo-Adjacency Set $\mathcal{A}$ is binary for $\mathcal{O}(\textsf{tw}^{2})$ pairs, contributing $2^{\mathcal{O}(\textsf{tw}^{2})}$ choices.
-Multiplying these components yields the total number of distinct signatures:
-$\text{Total States}=2^{\mathcal{O}(\textsf{tw}\log\textsf{tw})}\cdot 2^{\mathcal{O}(\textsf{tw}\log\textsf{tw})}\cdot 2^{\mathcal{O}(\textsf{tw}^{2})}\cdot 2^{\mathcal{O}(\textsf{tw}^{2})}=2^{\mathcal{O}(\textsf{tw}^{2})}$
-The DP algorithm processes each of the $\mathcal{O}(|V(G)|)$ nodes in the tree decomposition.
-At Leaf, Introduce, and Forget Nodes, the transition involves iterating over the $2^{\mathcal{O}(\textsf{tw}^{2})}$ signatures of the single child node, with the compatibility checks and updates for each taking time polynomial in the bag size, $\mathcal{O}(\textsf{tw}^{\mathcal{O}(1)})$.
-At Join Nodes, the transition requires checking all pairs of compatible signatures $(S_{1},S_{2})$ from the two child nodes.
-Since there are $2^{\mathcal{O}(\textsf{tw}^{2})}$ signatures at each child, there are at most $\left(2^{\mathcal{O}(\textsf{tw}^{2})}\right)^{2}=2^{\mathcal{O}(\textsf{tw}^{2})}$ pairs to examine, and the combining operations ($\oplus$ and $\otimes$) for each pair take time $\mathcal{O}(\textsf{tw}^{\mathcal{O}(1)})$.
-Consequently, the time required to process any single node is bounded by $2^{\mathcal{O}(\textsf{tw}^{2})}\cdot\textsf{tw}^{\mathcal{O}(1)}=2^{\mathcal{O}(\textsf{tw}^{2})}$.
-The overall running time of the algorithm is the number of nodes in the decomposition multiplied by the time per node:
-$\text{Time}=\mathcal{O}(|V(G)|)\cdot 2^{\mathcal{O}(\textsf{tw}^{2})}$
-This is bounded by $2^{\mathcal{O}(\textsf{tw}^{2})}\cdot|V(G)|^{\mathcal{O}(1)}$. This concludes the proof of the algorithmic result mentioned in Theorem 1. ∎
+(Algorithm in Theorem 1) The algorithm employs dynamic programming over a nice tree decomposition of $G\cup H$, maintaining states indexed by signatures $S=(\mathcal{P},\tau,\mathcal{A}_{t},\mathcal{R}_{t})$.
+For a bag $X_{t}$ of size at most $\textsf{tw}+1$, the total number of distinct signatures $|\mathcal{S}|$ is bounded by combining the state space components.
+The number of partitions $\mathcal{P}$ of $X_{t}$ is bounded by $2^{\mathcal{O}(\textsf{tw}\log\textsf{tw})}$.
+The status function $\tau$, which maps each part to one of $\mathcal{O}(\textsf{tw})$ possible representatives or status markers ($\uparrow,\downarrow$), yields at most $(\textsf{tw}+3)^{|\mathcal{P}|}=2^{\mathcal{O}(\textsf{tw}\log\textsf{tw})}$ possibilities.
+Since the number of pairs of parts is $\mathcal{O}(\textsf{tw}^{2})$, the total number of combinations for the adjacency sets $\mathcal{A}_{t}$ and $\mathcal{R}_{t}$ is $2^{\mathcal{O}(\textsf{tw}^{2})}$.
+Combining these, the total number of signatures per node is $|\mathcal{S}|\leq 2^{\mathcal{O}(\textsf{tw}^{2})}$.
+Each DP transition at Leaf, Introduce, and Forget nodes iterates over $\mathcal{O}(|\mathcal{S}|)$ states of the child, performing checks in polynomial time in tw.
+At Join nodes, pairs of states are combined, resulting in $\mathcal{O}(|\mathcal{S}|^{2})=2^{\mathcal{O}(\textsf{tw}^{2})}$ combined states, merged in polynomial time.
+Since the tree decomposition has $\mathcal{O}(|V(G)|)$ nodes, the overall running time is $\mathcal{O}(|V(G)|)\cdot 2^{\mathcal{O}(\textsf{tw}^{2})}=2^{\mathcal{O}(\textsf{tw}^{2})}\cdot|V(G)|^{\mathcal{O}(1)}$.
+This complexity proves the correctness of the algorithm mentioned in Theorem 1.
+∎
 3.2 Conditional Lower Bound
-In this section, we demonstrate that unless (ETH) fails, Labeled Contraction does not admit an algorithm running in time $2^{o(\textsf{tw}^{2})}\cdot|V(G)|^{\mathcal{O}(1)}$, where tw is the treewidth of the input graph $G$.
+In this section, we demonstrate that unless the ETH fails, Labeled Contraction does not admit an algorithm running in time $2^{o(\textsf{tw}^{2})}\cdot|V(G)|^{\mathcal{O}(1)}$, where tw is the treewidth of the input graph $G$.
 To obtain the lower bound, we reduce from a special case of Vertex Cover on sub-cubic graphs, defined in [3] as follows:
@@
 We prove Theorem 3 and Theorem 4 in the following two subsections.
-See 3 See 4
 5.1 Conditional Lower Bound
@@
 We now reduce 11-in-33-SAT to Labeled Contraction, ensuring that the number of vertices in $G$ and $H$ is linearly bounded by the number of variables and clauses in and instance of 11-in-33-SAT.
-We now proceed with the proof of Theorem 3.
+We now proceed with the proof of Theorem 3 which we restate here.
+See 3
 Proof.
@@
 5.2 Algorithmic Result
-We first present proof of Theorem 4 establishing an alternate $\mathsf{FPT}$ algorithm and later mention its consequences.
-Recall that the theorem states that the Labeled Contraction problem admits an algorithm with running time $(\delta(H)+1)^{k}\cdot|V(G)|^{\mathcal{O}(1)}$, where $k=|V(G)\setminus V(H)|$.
+We first present proof of Theorem 4 establishing an alternate $\mathsf{FPT}$ algorithm and later mention its consequences. We restate the theorem.
+See 4
+Proof.
 Let $(G,H)$ be an instance of the Labeled Contraction problem. With sanity check, we can assume that $V(H)\subseteq V(G)$.
@@
 6 Brute-force Algorithm and its Optimality
-We now prove Theorem 5. We restate it for reader’s conveience.
+We now prove Theorem 5. We restate it for reader’s convenience.
 See 5
+6.1 Brute-force Algorithm
+We prove the following lemma.
 Lemma 6.1.
@@
 ∎
-Now we present the conditional lower bound, demonstrating that the algorithm is essentially optimal unless the Exponential Time Hypothesis (ETH) fails.
+6.2 Conditional Lower Bound
+In this subsection, we present the conditional lower bound, demonstrating that the algorithm is essentially optimal unless the Exponential Time Hypothesis (ETH) fails.
 Our lower bound is established through a polynomial-time reduction from the Cross Matching problem, introduced by Fomin et al.